### v1

# The Euler class of the normal bundle of a Seifert fibration and horizontal foliations

Abstract.
For Seifert fibred manifolds with orientable base orbifolds, we establish a necessary and sufficient condition for the Euler class of the normal bundle of the Seifert fibration to vanish. When the base orbifold is hyperbolic, we also provide a second proof of this condition from the perspective of discrete faithful representations of Fuchsian groups.
As an application, we present infinitely many Seifert fibred rational homology spheres that admit co-oriented taut foliations but none with vanishing Euler class. In the context of the $L$-space conjecture, these provide examples of rational homology spheres that admit co-oriented taut foliations (and hence are not $L$-spaces) and have left-orderable fundamental groups yet none of the left-orders arise directly from the universal circle actions associated to co-oriented taut foliations.
Steven Boyer was partially supported by NSERC grant RGPIN 2024-06196
Ying Hu was partially supported by NSF grant DMS-2409398
2010 Mathematics Subject Classification. Primary 57M25, 57M50, 57M99
Key words: Seifert fibred manifold, normal bundle, Euler class, horizontal foliation
1. Introduction
In this article, we study the triviality of the normal bundle $\nu_{M}$ to the Seifert fibres of a Seifert fibration on a $3$-manifold $M$ with orientable base orbifold. We will refer to $\nu_{M}$ as the normal bundle of the Seifert fibration. As an oriented $2$-plane bundle, $\nu_{M}$ is trivial if and only if its Euler class $e(\nu_{M})\in H^{2}(M)$ vanishes.
We denote a closed oriented Seifert fibred manifold $M$ with orientable base orbifold by $M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, where $b\in\mathbb{Z}$ and $b_{i}\geq 2$ for $i=1,\ldots,n$. See §2.1 for a detailed explanation of the notation. We denote the base orbifold by $\mathcal{B}$; this is a closed orientable surface $|\mathcal{B}|$ of genus $g\geq 0$ with $n\geq 0$ cone points of orders $b_{1},...,b_{n}$. Recall that the Euler characteristic $\chi(\mathcal{B})$ of $\mathcal{B}$ is $\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$, and the Euler number of $M$ is $e(M)=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}$.
Here is our main result.
Theorem 1.1.
Given a closed oriented Seifert fibred manifold $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes if and only if there exists $m\in\mathbb{Z}$ such that
  1. (1)
    $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n,\;\text{and}$
  2. (2)
    $me(M)=\chi(\mathcal{B})$.
Remark 1.2.
It is often easy to identify an initial $m\in\mathbb{Z}$ that satisfies the congruence equations $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i})$ for $i=1,\cdots,n$. Any other integer $m^{\prime}$ satisfying these congruences can be written as $m^{\prime}=m+rl$, where $l$ is the least common multiple of $b_{1},b_{2},\ldots,b_{n}$ and $r\in\mathbb{Z}$. Hence when $n>0$, $e(M)\neq 0$ and $\chi(\mathcal{B})\neq 0$, the condition $m\,e(M)=\chi(\mathcal{B})$ in Theorem 1.1 can be replaced by
$\frac{\chi(\mathcal{B})}{e(M)}\text{ is an integer congruent to }m\;(\operatorname{mod}l)$
Next, we deduce the following easily verifiable necessary condition for $\nu_{M}$ to be a trivial bundle:
Corollary 1.3.
Given a closed oriented Seifert fibred manifold $M$, if the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes, then either $M$ is an oriented $S^{1}$-bundle over the torus or one of the following holds:
  1. (1)
    $e(M)=\chi(\mathcal{B})=0$,
  2. (2)
    $e(M)\neq 0,\,\chi(\mathcal{B})\neq 0,$ and
    $\frac{\chi(\mathcal{B})}{e(M)}\in\mathbb{Z}.$
Note that if $M$ is an oriented $S^{1}$-bundle over the torus, then its normal bundle is always trivial (see Remark 2.2).
By abelianizing the presentation (3.0.1) of $\pi_{1}(M)$ obtained by viewing $M$ as doing a Dehn filling of a trivial $S^{1}$-bundle over an $n$-punctured orientable surface of genus $g$ as in §2.1, one can show that if $T_{1}(M)$ denotes the torsion subgroup of $H_{1}(M)$, then
$|T_{1}(M)|=b_{1}\cdots b_{n}|e(M)|,$
where we take the convention that $b_{1}b_{2}\cdots b_{n}=1$ when $n=0$. Thus, it follows from Corollary 1.3 that:
Corollary 1.4.
Consider the Seifert fibred manifold $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ and suppose that the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes. Then the integer $b_{1}\ldots b_{n}\chi(\mathcal{B})$ is a multiple of $|T_{1}(M)|$. Hence, if $M$ is a rational homology $3$-sphere, then $b_{1}\ldots b_{n}\chi(\mathcal{B})$ is a multiple of $|H_{1}(M)|$.
Our initial motivation for studying the Euler class of $\nu_{M}$ comes from the study of horizontal foliations on Seifert fibred manifolds. By definition, a horizontal foliation is transverse to the Seifert fibration and as such, their tangent bundles are isomorphic to $\nu_{M}$. Hence, their Euler classes coincide, at least up to sign.
It is predicted by the $L$-space conjecture [3, 13] that if a closed, orientable, prime $3$-manifold admits a co-orientable taut foliation, then its fundamental group is left-orderable. One of the main tools for proving left-orderability of the fundamental group in the presence of a co-orientable taut foliation is to use Thurston’s universal circle action. See [7, 4] for the details. However, this approach requires the Euler class of the taut foliation to vanish. Although it was expected that examples of $3$-manifolds admitting co-orientable taut foliations but none with vanishing Euler class should exist, no published examples were previously known. In this article, we settle the matter by presenting such examples within the family of Seifert fibred rational homology $3$-spheres. Our arguments are based on the following application of Theorem 1.1.
Theorem 1.5.
Suppose that $M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ is a rational homology $3$-sphere which admits a co-orientable taut foliation $\mathcal{F}$. Then $\mathcal{F}$ has zero Euler class if and only if there exists $m\in\mathbb{Z}$ such that
  1. (1)
    $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n,\;\text{and}$
  2. (2)
    $me(M)=\chi(\mathcal{B})$.
The simplest examples of manifolds in which every co-oriented taut foliation has non-zero Euler class arise from Dehn surgery on the trefoil knot. In the following statement, we assume that $p$ and $q$ are nonzero coprime integers.
Theorem 1.6.
The $3$-manifold obtained by $p/q$-surgery along the right-handed trefoil knot admits a co-oriented taut foliation if and only if $p/q<1$. It admits one with zero Euler class if and only if $p/q<1$ and $|q|\equiv 1\pmod{p}$.
We will present more examples of rational homology Seifert fibred manifolds that admit co-oriented taut foliations, but none with vanishing Euler class, in §4.3 (Examples 4.4). In fact, our computations in §4 suggests that most closed Seifert fibred manifolds have nontrivial normal bundle, including the case when the base orbifold is Euclidean (i.e. $\chi(\mathcal{B})=0$). This may seem counterintuitive, since we have pointed out that the normal bundle is always trivial when the base $\mathcal{B}$ is the torus.
We close with an interesting by-product of our analysis.
Theorem 1.7.
Let $M$ be a closed oriented Seifert fibred manifold with orientable base orbifold $\mathcal{B}$. If $\mathcal{B}$ is hyperbolic or Euclidean, and $e(\nu_{M})=0$, then $M$ admits a horizontal foliation.
In contrast, closed Seifert fibred manifolds with spherical base orbifolds have finite fundamental groups and hence cannot admit taut foliations. Yet, in §4.1, we will show examples of such manifolds for which $e(\nu_{M})$ vanishes.
Plan of the paper
In Section 2 we explain our notation for Seifert manifolds and prove Theorem 1.1, which gives necessary and sufficient conditions for the Euler class of the normal bundle of a Seifert fibration to vanish. We also prove Corollary 1.3 in this section. In Section 3 we present a different perspective on Theorem 1.1 and Corollary 1.3 in the case that the base orbifold is hyperbolic through the use of discrete faithful $PSL(2,\mathbb{R})$-representations of $\pi_{1}(\mathcal{B})$. As a consequence of this approach, we prove the hyperbolic case of Theorem 1.7. Finally, in Section 4 we prove Theorem 1.5 and Theorem 1.6 and provide examples of Seifert fibred manifolds whose normal bundles are (non-)trivial. We classify which occurs when $\mathcal{B}$ is spherical in §4.1 and is Euclidean in §4.2. The Euclidean case of Theorem 1.7 is dealt with at the end of §4.2. The case that $\mathcal{B}$ is hyperbolic is analysed in §4.3. Finally, in §4.4, we illustrate how the conditions of Theorem 1.1 can be used to verify the known fact that $e(\nu_{M})=0$ when $|H_{1}(M)|=1,2$.
Acknowledgements
We thank Nathan Dunfield for bringing up the question of finding examples of rational homology spheres that admit co-oriented taut foliations but none with vanishing Euler class, during his lecture at the workshop “Low Dimensional Topology and Floer Theory” in Montreal, August 18-29, 2025, which inspired us to investigate the Euler class of horizontal foliations in Seifert fibred manifolds. The second and third named authors would like to thank the Simons Foundation, the CRM, and CIRGET for their support during the preparation of this article.
2. The Euler class of the normal bundle of a Seifert fibration
Let $M$ be a closed, oriented Seifert fibred manifold whose base orbifold has an orientable underlying space. In this section, we compute the Euler class of the normal bundle of the Seifert fibration on $M$ and determine conditions for it to vanish (Theorem 1.1 and Corollary 1.3).
2.1. Seifert notation
We first recall the Seifert notation that we will use to describe $M$.
Consider $M$ constructed by attaching $n+1$ solid tori $N_{i}$, $i=0,1,\dots,n$, to a trivial $S^{1}$-bundle over an orientable surface $F$ with $|\partial F|=n+1$, where we assume that $N_{0}$ is a tubular neighbourhood of a regular fibre and the $N_{i}$ are tubular neighbourhoods of the singular fibres for $i=1,2,\ldots,n$. Denote this trivial $S^{1}$-bundle by $P\cong F\times S^{1}$. Fix orientations on $F$ and on the fibres so that they intersect positively.
Let
$\partial F=\mu_{0}\sqcup\mu_{1}\sqcup\cdots\sqcup\mu_{n}$
with the induced orientation, and
$\partial P=T_{0}\sqcup T_{1}\sqcup\cdots\sqcup T_{n}$
with $\mu_{i}\subset T_{i}$ for each $i=0,1,\cdots,n$. So on each $T_{i}$, we use the basis given by $\{h_{i},\mu_{i}\}$, where $h_{i}$ is given by a Seifert fibre. By our orientation convention, we have $\mu_{i}\cdot h_{i}=1$.
Let $D_{i}$ be a meridional disk of $N_{i}$, oriented so that it intersects the Seifert fibres positively. Under the attaching map, we have $\partial D_{i}=-\bigl(a_{i}h_{i}+b_{i}\mu_{i}\bigr)$ as oriented curves, for some $a_{i}\in\mathbb{Z}$, $b_{i}\in\mathbb{Z}_{>0}$ and $a_{i}$, $b_{i}$ coprime. In fact, since $N_{0}$ is a tubular neighbourhood of a regular fibre, and the $N_{i}$ are tubular neighbourhoods of singular fibres, we have $b_{0}=1$ and $b_{i}\geq 2$.
Let $g\geq 0$ denote the genus of $F$, $b=\frac{a_{0}}{b_{0}}\in\mathbb{Z}$, and denote
$M=M\Bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\Bigr)$
Let $\mathcal{B}$ denote the base orbifold of $M$, which is a genus $g$ closed orientable surface with $n$ cone points of orders $b_{1},\ldots,b_{n}$ respectively. We use $|\mathcal{B}|$ to denote the underlying space of $\mathcal{B}$.
Remarks 2.1.
  1. (1)
    This description of the Seifert fibred manifold is not unique. More precisely, we have that $M\Bigl(g;\,\tfrac{a_{0}}{b_{0}},\ldots,\tfrac{a_{n}}{b_{n}}\Bigr)$ is homeomorphic to $M\Bigl(g;\,\tfrac{a^{\prime}_{0}}{b_{0}},\ldots,\tfrac{a^{\prime}_{n}}{b_{n}}\Bigr)$ by an orientation-preserving homeomorphism preserving fibres if and only if up to reordering,
    $\frac{a_{i}}{b_{i}}-\frac{a^{\prime}_{i}}{b_{i}}\in\mathbb{Z}\quad\text{and}\quad\sum_{i=0}^{n}\frac{a_{i}}{b_{i}}=\sum_{i=0}^{n}\frac{a^{\prime}_{i}}{b_{i}}.$
    By choosing exactly one of the $N_{i}$, namely $N_{0}$, to be a neighbourhood of a regular fibre in the construction above, we are adopting the usual convention of having only one of the $\tfrac{a_{i}}{b_{i}}$ be an integer.
  2. (2)
    Given $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, the sum
    $e(M):=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}$
    is an invariant of the given Seifert fibring of $M$ called the Euler number.
Remark 2.2.
When $n=0$, the manifold $M=M(g;b)$ is an $S^{1}$-bundle over a closed, orientable surface $B$ of genus $g$ with Euler number $b$. Moreover, the necessary and sufficient condition in Theorem 1.1 simplifies to the existence of $m\in\mathbb{Z}$ such that
(2.1.1) $mb=\chi(B)$
By assumption, the Euler class of the circle bundle $f:M\to B$ is $bu$, where $u$ is a generator of $H^{2}(B)\cong\mathbb{Z}$. It is easy to see that $\nu_{M}$ is the pull-back by $f$ of the tangent bundle $\tau_{B}$ of $B$. Hence,
(2.1.2) $e(\nu_{M})=f^{*}(e(\tau_{B})),$
Let $\langle h\rangle$ denote the central subgroup of $\pi_{1}(M)$ generated by a circle fibre of $f$ and consider the $5$-term exact sequence in cohomology associated to the central extension $1\to\langle h\rangle\to\pi_{1}(M)\to\pi_{1}(B)\to 1$:
$0\to H^{1}(B)\xrightarrow{}H^{1}(M)\to\mathbb{Z}\xrightarrow{\;\delta\;}H^{2}(B)\xrightarrow{\;f^{*}\;}H^{2}(M)$
It follows from [8, Theorem 4] that the image of 1-1 under the transgression homomorphism $\delta$ is the Euler class $bu$ of the circle bundle. Hence $\mbox{kernel}(f^{*})$ is generated by $bu$. Then by (2.1.2), $e(\nu_{M})=0$ if and only $e(\tau_{B})=\pm\chi(B)u$ is a multiple of $b$, which is condition (2.1.1).
We consider the following two special cases:
  1. (1)
    If $M=M(g;b)$ is a rational homology sphere, then the base $B$ must be the $2$-sphere, and $M=M(0;b)$ is a lens space $L(|b|,1)$. In this case, the Euler class $e(\nu_{M})$ vanishes if and only if $b\in\{\pm 1,\pm 2\}$.
  2. (2)
    When $B$ is the torus, (2.1.2) implies that the Euler class $e(\nu_{M})$ is always trivial, since the tangent bundle $T\mathcal{B}$ is trivial,
2.2. Proof of Corollary 1.3
We prove Corollary 1.3 using Theorem 1.1 here and defer the proof of Theorem 1.1 to the next subsection.
Let $\nu_{M}$ denote the normal plane bundle over $M$ to the Seifert fibration, oriented with the induced orientation from those on $M$ and on the fibres.
Proof of Corollary 1.3.
Assume that the Euler class of the normal bundle vanishes.
We first consider the case $n=0$, in which $\mathcal{B}$ is a genus $g$ surface and $M$ is an $S^{1}$-bundle over $\mathcal{B}$ of Euler number $b$. The necessary and sufficient condition in Theorem 1.1 is equivalent to that the existence of $m\in\mathbb{Z}$ such that $mb=\chi(\mathcal{B})$.
If $m=0$, then $\chi(\mathcal{B})=mb=0$ and therefore $M$ is an oriented $S^{1}$-bundle over the torus. On the other hand, if $M$ is not an oriented $S^{1}$-bundle over the torus, then $\chi(\mathcal{B})\neq 0$. Then for the equality $mb=\chi(\mathcal{B})$ to hold, we must have $e(M)=b\neq 0$. Hence
$m=\frac{\chi(\mathcal{B})}{e(M)}\in\mathbb{Z}.$
Next we consider the case $n>0$. Again by Theorem 1.1, there exists $m\in\mathbb{Z}$ such that $ma_{i}\equiv 1\pmod{b_{i}}$ and $me(M)=\chi(\mathcal{B})$. Since $m$ cannot be zero in this case, either (1) or (2) holds. ∎
2.3. Proof of Theorem 1.1
Here we provide a proof of Theorem 1.1. When the base orbifold is hyperbolic, i.e., $\chi(\mathcal{B})<0$, we will give another perspective on Theorem 1.1 and Corollary 1.3 in §3. There we will show that the Euler class of the normal bundle is precisely the pullback of the Euler class of a discrete faithful representation of $\pi_{1}(\mathcal{B})$ in $PSL(2,\mathbb{R})$ under the homomorphism $\pi_{1}(M)\to\pi_{1}(\mathcal{B})$ induced by the natural projection $M\rightarrow\mathcal{B}$  [^1] Since $M$ is irreducible when $\mathcal{B}$ is hyperbolic, we have $H^{2}(M)\equiv H^{2}(\pi_{1}(M))$..
Proof of Theorem 1.1.
To simplify the notation in the proof, we will use $\tfrac{a_{0}}{b_{0}}$ in place of $b$, as explained in §2.1. Thus
$M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)=M\bigl(0;\,\tfrac{a_{0}}{b_{0}},\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$
We will compute the Euler class of the normal bundle, denoted by $e(\nu_{M})\in H^{2}(M)$, using the following exact sequence of the pair $(M,\sqcup_{i}T_{i})$:
$\cdots\to\oplus_{i}H^{1}(T_{i})\xrightarrow{\delta}H^{2}(M,\sqcup_{i}T_{i})\xrightarrow{i^{*}}H^{2}(M)\to 0$
Note that
$H^{2}(M,\sqcup_{i}T_{i})\cong H^{2}(P,\sqcup_{i}T_{i})\oplus H^{2}(N_{1},T_{1})\oplus\cdots\oplus H^{2}(N_{n},T_{n}),$
and by the Künneth formula,
$H^{2}(P,\sqcup_{i}T_{i})\cong H^{1}(F,\partial F)\otimes H^{1}(S^{1})\oplus H^{2}(F,\partial F)$
Recall that $\partial F=\sqcup_{i}\mu_{i}$ with $\mu_{i}\subset T_{i}$ for each $i=0,1,\ldots,n$. Suppose that oriented simple closed curves $\alpha_{r},\beta_{r}$, $r=1,\ldots,g$, together with oriented arcs $c_{j}$ on $F$ that connect the boundary component $\mu_{j}$ to the boundary component $\mu_{j+1}$, $j=0,1,\ldots,n-1$ when $n>0$, represent a set of generators of $H_{1}(F,\partial F)$. Then $H_{2}(P,\sqcup_{i}T_{i})$ is freely generated by the classes of vertical tori $\alpha_{r}\times S^{1}$, $\beta_{r}\times S^{1}$, the horizontal surface $F$ and vertical annuli $c_{j}\times S^{1}$ when $n>0$. We denote their duals in $H^{2}(P,\partial P)$ by
$\alpha_{r}^{*}\otimes S^{*},\quad\beta_{r}^{*}\otimes S^{*},\quad c_{j}^{*}\otimes S^{*},\quad F^{\ast}$
respectively, where $r=1,\ldots,g$ and $j=0,1,\ldots,n-1$.
Finally, let $D_{i}^{*}\in H^{2}(N_{i},T_{i})$ denote the dual of $D_{i},\partial D_{i}]\in H_{2}(N_{i},T_{i})$ for each $i=0,1,\ldots,n$.
In summary, a cohomology class in
$H^{2}(M,\sqcup T_{i})\cong H^{1}(F,\partial F)\otimes H^{1}(S^{1})\oplus H^{2}(F,\partial F)\oplus H^{2}(N_{1},T_{1})\oplus\cdots\oplus H^{2}(N_{n},T_{n})$
can be uniquely expressed as a linear combination of
$\alpha_{r}^{*}\otimes S^{*},\quad\beta_{r}^{*}\otimes S^{*},\quad F^{*},\quad D_{i}^{*}\quad\text{and}\quad c_{j}^{*}\otimes S^{*}\text{ when }n>0$
where $r=1,\ldots,g$, $j=0,1,\ldots,n-1$ and $i=0,1,\ldots,n$.
The proof proceeds in two steps (1) and (2).
  1. (1)
    Find the preimage of the Euler class $e(\nu_{M})$ of the normal bundle $\nu_{M}$ in $H^{2}(M,\sqcup T_{i})$.
    Let $\sigma$ be a nowhere vanishing vector field along $\partial P=\sqcup T_{i}$ that points out of $P$. Since $T_{i}$ is tangent to the Seifert fibres, $\sigma$ is a section of $\nu_{M}$ along $\sqcup T_{i}$. We give a cellular structure on $M$ for which we assume that the vertical surfaces $\alpha_{r}\times S^{1}$, $\beta_{r}\times S^{1}$, $c_{j}\times S^{1}$, and the horizontal surfaces $F$ and $D_{i}$, as well as the $T_{i}$, all belong to the 2-skeleton. We also assume that the curves $\alpha_{r}$, $\beta_{r}$ and the arcs $c_{j}$ on $F$ are in the 1-skeleton.
    From $\sigma$ one can define an obstruction $2$-cochain, denoted by $c_{\sigma}$, that represents the relative Euler class of the normal bundle $\nu_{M}$ in $H^{2}(M,\sqcup T_{i})$. By construction, the image of $c_{\sigma}]$ under $H^{2}(M,\sqcup T_{i})\to H^{2}(M)$ is the Euler class $e(\nu_{M})$. (See [9, Section 2] for instance.)
    To write $c_{\sigma}]$ as a linear combination of $\alpha_{r}^{*}\otimes S^{*}$, $\beta_{r}^{*}\otimes S^{*}$, $c_{j}^{*}\otimes S^{*}$, $F^{*}$, and $D_{i}^{*}$, it suffices to compute its evaluation on the dual homology classes.
    1. (a)
      $c_{\sigma}]$ evaluates to zero on all vertical surfaces. This is because, starting from $\sigma|_{\partial F}$, where $\partial F\subset\partial P=\sqcup T_{i}$, there is no obstruction to extending $\sigma$ to the 1-skeleton of $F$. Since $P\cong F\times S^{1}$ is a trivial $S^{1}$-bundle, one can extend the section further to the vertical surfaces $\alpha_{r}\times S^{1}$, $\beta_{r}\times S^{1}$ and $c_{j}\times S^{1}$ by staying constant along the fibre direction.
    2. (b)
      By the Poincaré-Hopf Theorem, $c_{\sigma}]$ evaluated on a horizontal surface equals its Euler characteristic, i.e.,
      $c_{\sigma}]([D_{i},\partial D_{i}])=1,\quad i=0,1,\ldots,n,$
      and
      $c_{\sigma}]([F,\partial F])=\chi(F)=\chi(|\mathcal{B}|)-(n+1).$
  2. (2)
    Compute the image of $\delta:\oplus_{i}H^{1}(T_{i})\rightarrow H^{2}(M,\sqcup_{i}T_{i})$.
    Let $h_{i}^{*},\mu_{i}^{*}\in H^{1}(T_{i})$ denote the (Hom) dual classes of $h_{i}],[\mu_{i}]\in H_{1}(T_{i})$ for $i=0,1,\ldots,n$. We compute $\delta h_{i}^{*}$ and $\delta\mu_{i}^{*}$, which generate the image of $\delta$.
    1. (a)
      It is easy to see that, since $\delta h_{i}^{*}=h_{i}^{*}\circ\partial$ and $\delta\mu_{i}^{*}=\mu_{i}^{*}\circ\partial$, they vanish on all vertical tori, $\alpha_{r}\times S^{1}$ and $\beta_{r}\times S^{1}$.
    2. (b)
      As $\partial([c_{j}\times S^{1}])=[h_{j+1}]-[h_{j}]$ in $H_{1}(\sqcup_{i}T_{i})$ for $j=0,1,\ldots,n-1$, we have
      $\delta\mu_{i}^{*}([c_{j}\times S^{1}])=0\quad\text{for all }i=0,1,\ldots,n,\text{ and }\ j=0,1,\ldots,n-1,$
      and
      $\delta h_{i}^{*}([c_{j}\times S^{1}])=\begin{cases}1&\text{if }j=i-1\\[6.0pt] -1&\text{if }j=i\\[6.0pt] 0&\text{otherwise}\end{cases}$
    3. (c)
      Since $\partial F=\sqcup_{i}\mu_{i}$, we have $\delta h_{i}^{*}([F,\partial F])=0$ and $\delta\mu_{i}^{*}([F,\partial F])=1$ for all $i=0,1,\ldots,n.$
    4. (d)
      Finally, to compute the values of $\delta h_{i}^{*}$ and $\delta\mu_{i}^{*}$ over $D_{i},\partial D_{i}]\in H_{2}(N_{i},T_{i})$, we note that $\partial[D_{i}]=-(a_{i}[h_{i}]+b_{i}[\mu_{i}])$ by our orientation convention in order to keep $b_{i}>0$ (see §2.1). Then it follows that for $i=0,1,\ldots,n$
      $\delta h_{i}^{*}([D_{i},\partial D_{i}])=h_{i}^{*}(-a_{i}[h_{i}]-b_{i}[\mu_{i}])=-a_{i}$
      and
      $\delta\mu_{i}^{*}([D_{i},\partial D_{i}])=\mu_{i}^{*}(-a_{i}[h_{i}]-b_{i}[\mu_{i}])=-b_{i}.$
In summary, we have
$c_{\sigma}]=\chi(F)F^{*}+\sum_{i=0}^{n}D_{i}^{*}$
The image of $\delta$ is generated by
$\delta\mu_{i}^{*}=F^{*}-\sum_{i=0}^{n}b_{i}D_{i}^{*},\quad\text{ and }\quad\delta h_{i}^{*}=\begin{cases}-c_{0}^{*}\otimes S^{*}-a_{0}D_{0}^{*},&\text{for }i=0,\\[6.0pt] c_{i-1}^{*}\otimes S^{*}-c_{i}^{*}\otimes S^{*}-a_{i}D_{i}^{*},&\text{for }0<i<n,\\[6.0pt] c_{n-1}^{*}\otimes S^{*}-a_{n}D_{n}^{*},&\text{for }i=n.\end{cases}$
if $n>0$; and when $n=0$, the image of $\delta$ is generated by
$\delta\mu_{0}^{*}=F^{*}-b_{0}D_{0}^{*}=F^{*}-D_{0}^{*}\quad\text{and}\quad\delta h_{0}^{*}=-a_{0}D_{0}^{*}=-bD_{0}^{*}.$
The Euler class $e(\nu_{M})=i^{*}([c_{\sigma}])\in H^{2}(M)$ vanishes if and only if $c_{\sigma}]\in\mathrm{Im}(\delta)$. That is, there exist integers $m_{i},n_{i}$, for $i=0,1,\ldots,n$, such that
$\delta\Bigl(\sum_{i}-m_{i}h_{i}^{*}+\sum_{i}n_{i}\mu_{i}^{*}\Bigr)=[c_{\sigma}].$
  1. (1)
    When $n>0$, for the coefficients of $c_{j}^{*}\otimes S^{*}$ to be zero, the $m_{i}$ must be identical. Hence it follows that $c_{\sigma}]\in\mathrm{Im}(\delta)$ if and only if there exist $m,n_{i}\in\mathbb{Z}$ for $i=0,\ldots,n$ such that
    $\begin{cases}ma_{i}-n_{i}b_{i}=1,\mbox{ for }i=0,1,\ldots,n\\[6.0pt] \sum_{i=0}^{n}n_{i}=\chi(F)=\chi(|\mathcal{B}|)-(n+1).\end{cases}$
    Recall that $\tfrac{a_{0}}{b_{0}}=b$. Hence the equation $ma_{0}-n_{0}b_{0}=1$ simplifies to $mb-n_{0}=1$, and therefore, the condition above is equivalent to the existence of $m,n_{i}\in\mathbb{Z}$ for $i=1,\ldots,n$ such that
    $\begin{cases}ma_{i}-n_{i}b_{i}=1,\text{ for }i=1,\ldots,n\\[6.0pt] mb-1+\displaystyle\sum_{i=1}^{n}n_{i}=\chi(F)=\chi(|\mathcal{B}|)-(n+1)\end{cases}$
    For $i\geq 1$, we have
    $n_{i}=\frac{ma_{i}-1}{b_{i}}=\frac{ma_{i}}{b_{i}}-\frac{1}{b_{i}}$
    Hence, the second condition is equivalent to
    $m(b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}})=\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$
    Note that
    $e(M)=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}\quad\text{and}\quad\chi(\mathcal{B})=\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$
    This completes the proof of the theorem when $n>0$.
  2. (2)
    If $n=0$ (and hence there is no singular fibre), the condition above reduces to the existence of $m\in\mathbb{Z}$ such that
    $mb=\chi(\mathcal{B})$
    From the discussion above, one can easily verify that this is a necessary and sufficient condition for the Euler class to vanish in this special case.
This proves the theorem. ∎
3. Discrete representations of Fuchsian groups and the normal bundle of a Seifert fibration
In this section we provide alternative proofs of Theorem 1.1 and Corollary 1.3 from the perspective of discrete faithful representations of Fuchsian groups. We normalise our notation for oriented Seifert manifolds $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ by further requiring that $\tfrac{a_{i}}{b_{i}}\in(0,1)$ for each $i=1,\ldots,n$. See Remark 2.1.
Figure 1. The generators $\bar{x}_{i}$ in the presentation of $\Gamma=\pi_{1}(\mathcal{B})$
The fundamental group of $M$ has a presentation
(3.0.1) $\hskip 28.45274pt\langle x_{1},x_{2},\ldots,x_{n},w_{1},w_{2},\ldots,w_{g},z_{1},z_{2},\ldots,z_{g},h\;|\;h\mbox{ central},x_{i}^{b_{i}}=h^{-a_{i}}\mbox{ for each }i,$
$\hskip 56.9055pt[w_{1},z_{1}][w_{2},z_{2}]\cdots[w_{g},z_{g}]x_{1}x_{2}\cdots x_{n}=h^{b}\rangle,$
and $\Gamma=\pi_{1}(\mathcal{B})$ has a presentation
(3.0.2) $\hskip 28.45274pt\langle\bar{x}_{1},\ldots,\bar{x}_{n},\bar{w}_{1},\ldots,\bar{w}_{g},\bar{z}_{1},\ldots,\bar{z}_{g}\;|\;\bar{x}_{i}^{b_{i}}=1,[\bar{w}_{1},\bar{z}_{1}]\cdots[\bar{w}_{g},\bar{z}_{g}]\bar{x}_{1}\cdots\bar{x}_{n}=1\rangle$
These two presentations are related via a well-known central extension:
(3.0.3) $E_{M}:1\to\mathbb{Z}\to\pi_{1}(M)\xrightarrow{\;\alpha\;}\Gamma\to 1,$
Here $h$ corresponds to the image in $\pi_{1}(M)$ of $1\in\mathbb{Z}$, $\alpha(x_{i})=\bar{x}_{i}$ for each $i$, $\alpha(w_{j})=\bar{w}_{j}$ for each $j$ and $\alpha(z_{k})=\bar{z}_{k}$ for each $k$.
3.1. The second cohomology group of $\Gamma$
Jankins and Neuman determined a canonical isomorphism between $H^{2}(\Gamma)$ and the abelian group having generators $A_{0},A_{1},\ldots,A_{n}$ and relations $A_{0}=b_{i}A_{i}$ for $\leq i\leq n [11$]. (The case that $g=0$ is discussed in [11] and the general case in [10].) Each element of $H^{2}(\Gamma)$ can be uniquely expressed in the form
$dA_{0}+c_{1}A_{1}+\cdots+c_{n}A_{n},$
where $d,c_{1},\ldots,c_{n}\in\mathbb{Z}$ and $1\leq c_{i}\leq b_{i}-1$ for each value of $i$.
The Euler class $e(E_{M})\in H^{2}(\Gamma)$ is the cohomology class which classifies the extension (3.0.3). In [11] it was shown to be
(3.1.1) $e(E_{M})=-(bA_{0}+a_{1}A_{1}+a_{2}A_{2}+\cdots+a_{n}A_{n})$
This class arises as the image of 1-1 under the transgression homomorphism $\delta$ from the $5$-term exact sequence in cohomology of $E_{M}$:
(3.1.2) $0\to H^{1}(\Gamma)\xrightarrow{}H^{1}(M)\to\mathbb{Z}\xrightarrow{\;\delta\;}H^{2}(\Gamma)\xrightarrow{\;\alpha^{*}\;}H^{2}(M)$
See [8, Theorem 4]. Thus,
(3.1.3) $\mbox{kernel}(\alpha^{*})=\langle e(E_{M})\rangle$
3.2. The normal bundle of the Seifert fibring of $M$ and discrete representations
Since $\mathcal{B}$ is hyperbolic, its universal cover can be identified with $\mathbb{H}^{2}$ and the associated group of deck transformations with a discrete subgroup of $\mbox{Isom}_{+}(\mathbb{H}^{2})=PSL_{2}(\mathbb{R})$. This yields a discrete faithful representation $\bar{\varphi}:\Gamma\to PSL_{2}(\mathbb{R})$.
There is a Seifert fibring $\widehat{M}\to\mathbb{H}^{2}$ lifting $M\to\mathcal{B}$ for which the following diagram commutes
$\widehat{M}=\mathbb{H}^{2}\times S^{1}$$\mathbb{H}^{2}$$M\mathcal{B}$
where $\widehat{M}$ as the total space of an $S^{1}$-bundle over the contractible $\mathbb{H}^{2}$ can be identified with $\mathbb{H}^{2}\times S^{1}$. Hence, we can identify the universal cover of $M$ with $\mathbb{H}^{2}\times\mathbb{R}$, where the vertical lines $\{x\}\times\mathbb{R}$ universally cover the Seifert fibres of $M$. Denote the normal bundle to these vertical lines by $\tilde{\nu}$.
The fundamental group of $M$ acts on $\mathbb{H}^{2}$ via the composition $\varphi=\bar{\varphi}\circ\alpha$, which gives rise to a diagonal action $\psi$ of $\pi_{1}(M)$ on $\widetilde{M}\times\mathbb{H}^{2}$:
$\psi(\gamma)(x,y)=(\gamma\cdot x,\varphi(\gamma)(y))$
This action is equivariant with respect to the projection $\widetilde{M}\times\mathbb{H}^{2}\to\widetilde{M}$, so taking quotients determines a locally-trivial $\mathbb{H}^{2}$-bundle
$M\times_{\psi}\mathbb{H}^{2}=(\widetilde{M}\times\mathbb{H}^{2})/\pi_{1}(M)\to M$
Fix a Riemannian metric on $M$ and pull it back to $\widetilde{M}$. Since the group of deck transformations of the cover $\widetilde{M}\to M$ acts as isometries of this metric, the associated exponential map $\mbox{exp}:T\widetilde{M}\to\widetilde{M}$ is $\pi_{1}(M)$-equivariant, as is the composition
$E(\tilde{\nu}_{M})\xrightarrow{\;\mbox{{\footnotesize exp}}\;}\widetilde{M}=\mathbb{H}^{2}\times\mathbb{R}\xrightarrow{\;\mbox{{\footnotesize projection}}\;}\mathbb{H}^{2}$
Now proceed as in the proof of [2, Lemma A.3], starting from its third paragraph, to show that the $\mathbb{H}^{2}$-bundle $M\times_{\psi}\mathbb{H}^{2}\to M$ is topologically equivalent to the normal bundle $\nu_{M}$ and conclude, as in [2, Proposition A4], that $e(\nu_{M})=e(\varphi)=e(\bar{\varphi}\circ\alpha)$. Thus
(3.2.1) $e(\nu_{M})=\alpha^{*}(e(\bar{\varphi}))$
It is well-known that $e(\bar{\varphi})$, respectively $e(\varphi)$, is the obstruction to lifting the representation $\bar{\varphi}:\Gamma\to PSL_{2}(\mathbb{R})$, respectively $\varphi:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$, to $\widetilde{PSL}_{2}(\mathbb{R})$ [14, Lemma 2].
An expression for $e(\bar{\varphi})$ in terms of the Jankins-Neumann generators $A_{0},A_{1},\ldots,A_{n}$ of $H^{2}(\Gamma)$ can be determined following the method of §3 of [11]. To explain it, recall the central extension $1\to\mathbb{Z}\to\widetilde{PSL}_{2}(\mathbb{R})\to PSL_{2}(\mathbb{R})\to 1$. Let $h\in\widetilde{PSL}_{2}(\mathbb{R})$ correspond to $1\in\mathbb{Z}$.
By construction, the generator $\bar{x}_{i}$ of the presentation (3.0.2) of $\Gamma$ is mapped to a conjugate of a rotation of angle $\frac{2\pi}{b_{i}}$ in $PSL_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{S}^{1})$ by $\bar{\varphi}$, so lifts to a conjugate of translation by $\frac{1}{b_{i}}$ in $\widetilde{PSL}_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{R})$. Hence, if
$1\to\mathbb{Z}\to\widetilde{\Gamma}\to\Gamma\to 1$
is the pull-back central extension determined by $\bar{\varphi}$, then $\widetilde{\Gamma}$ has a presentation
$\langle\tilde{x}_{1},\ldots,\tilde{x}_{n},\tilde{w}_{1},\ldots,\tilde{w}_{g},\tilde{z}_{1},\ldots,\tilde{z}_{g}\;|\;\tilde{x}_{i}^{b_{i}}=h,[\tilde{w}_{1},\tilde{z}_{1}]\cdots[w_{g},\tilde{z}_{g}]\tilde{x}_{1}\cdots\tilde{x}_{n}=h^{-b}\rangle$
It is a classic result in the theory of Fuchsian groups that under our assumptions, $b=\chi(|\mathcal{B}|)-n$. As such, the discussion in §3 of [11] implies that
(3.2.2) $e(\bar{\varphi})=(n-\chi(|\mathcal{B}|))A_{0}-A_{1}-A_{2}-\cdots-A_{n}$
Proof of Theorem 1.1 when $\mathcal{B}$ is hyperbolic.
As $\nu_{M}$ is an oriented plane bundle, it is trivial if and only if $e(\nu_{M})=0$. Combining (3.2.1), (3.1.3), (3.1.1), and (3.2.2) shows that $e(\nu_{M})=0$ if and only if there is an integer $m$ such that
(3.2.3) $(\chi(|\mathcal{B}|)-n)A_{0}+A_{1}+\cdots+A_{n}=m(bA_{0}+a_{1}A_{1}+a_{2}A_{2}+\cdots+a_{n}A_{n})$
Consideration of the normal form for elements of $H^{2}(\Gamma)$ described in §3.1 shows that the existence of such an $m$ is equivalent to the existence of an integer $m$ for which the following two conditions hold
  1. (1)
    $ma_{i}\equiv 1\mbox{ (mod $b_{i}$) for }1\leq i\leq n$;
  2. (2)
    $mb+\sum_{i=1}^{n}\left\lfloor\frac{ma_{i}}{b_{i}}\right\rfloor=\chi(|\mathcal{B}|)-n$.
For each $i\in\{1,2,\ldots,n\}$, it follows from the congruence equation $ma_{i}\equiv 1\pmod{b_{i}}$ that
$\left\lfloor\frac{ma_{i}}{b_{i}}\right\rfloor=\frac{ma_{i}}{b_{i}}-\frac{1}{b_{i}}$
Hence, the second equation becomes
$m\left(b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}\right)=\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$
which again simplifies to
$me(M)=\chi(\mathcal{B})$
This completes the proof. ∎
Though we proved Corollary 1.3 as a consequence of Theorem 1.1 in §2.2, we provide a proof of the hyperbolic case of the corollary next which is independent of that theorem.
Proof of Corollary 1.3 when $\mathcal{B}$ is hyperbolic.
Let $l$ denote the least common multiple of the integers $b_{1},b_{2},\ldots,b_{n}$ and note that as $\gcd(\frac{l}{b_{1}},\frac{l}{b_{2}},\ldots,\frac{l}{b_{n}})=1$, we have an epimorphism
$\theta:H^{2}(\Gamma)\to\mathbb{Z},\;A_{0}\mapsto l,A_{i}\mapsto\frac{l}{b_{i}}\mbox{ for }1\leq i\leq n$
Then by (3.1.1),
(3.2.4) $\theta(e(E_{M}))=-l(b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}})=-le(M),$
where $e(M)$ is the Euler number of $M$ (see §2.1).
Next note that (3.2.2) implies that
(3.2.5) $\theta(e(\bar{\varphi}))=l(n-\chi(|\mathcal{B}|)-\sum_{i}\frac{1}{b_{i}})=-l\chi(\mathcal{B})$
Suppose that $e(\nu_{M})=0$. Then (3.2.1) and (3.1.3) imply that there is an integer $m$ such that $e(\bar{\varphi})=me(E_{M})\in H^{2}(\Gamma)$ and therefore $\theta(e(\bar{\varphi}))=m\theta(e(E_{M}))$. Since $\mathcal{B}$ is hyperbolic, $\chi(\mathcal{B})\neq 0$, and therefore (3.2.4) and (3.2.5) imply that $m\neq 0$ and $\chi(\mathcal{B})=me(M)$, which is what was to be proved. ∎
Proof of Theorem 1.7 when $\mathcal{B}$ is hyperbolic.
We saw in (3.2.1) that that $e(\nu_{M})=\alpha^{*}(e(\bar{\varphi}))=e(\bar{\varphi}\circ\alpha)=e(\varphi)$, where $\varphi=\bar{\varphi}\circ\alpha:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$. Hence, if $e(\nu_{M})=0$, then $\varphi$ lifts to a homomorphism $\pi_{1}(M)\to\widetilde{PSL}_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{R})$ [14, Lemma 2]. It follows that $\pi_{1}(M)$ is left-orderable [6, Theorem 1.1] and therefore, $M$ admits a horizontal foliation [6, Theorem 1.3]. ∎
4. Examples and applications
Recall that if $M$ is a Seifert fibred rational homology $3$-sphere, then the underlying surface of the base orbifold is either $S^{2}$ or $P^{2}$. However, in the latter case, $M$ does not admit a co-orientable taut foliation. (See [6] for instance.)
In this section we use Theorem 1.1 and Corollary 1.3 to analyse when an oriented Seifert manifold $M=M(0;b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$ has a trivial normal bundle and apply this to show that generically, if $M$ has a co-oriented taut foliation, it has none with zero Euler class.
We assume that $n\geq 3$. Then the base orbifold $\mathcal{B}$ of $M$ is spherical if $\chi(\mathcal{B})>0$, Euclidean if $\chi(\mathcal{B})=0$, and hyperbolic if $\chi(\mathcal{B})<0$. Orient $\mathcal{B}$ and endow $\nu_{M}$, the normal bundle to the Seifert fibres of $M$, with the induced orientation. The following lemma is key to our applications.
Lemma 4.1.
Suppose that $M$ is a Seifert fibred rational homology $3$-sphere. Then
$(1)$ Any co-orientable taut foliation on $M$ is horizontal. That is, it can be isotoped to be everywhere transverse to the Seifert fibres of $M$.
$(2)$ The tangent space to any co-orientable taut foliation on $M$ is isomorphic to $\nu_{M}$ as an unoriented bundle. In particular, the Euler class of any co-oriented taut foliation on $M$ agrees, up to sign, with that of $\nu_{M}$.
Proof.
That any co-oriented taut foliation on $M$ must be horizontal is [5, Corollary 7] in the case $n=3$ and [1, Proposition 6.1] in general. Thus (1) holds. Consequently, the tangent bundle of any such foliation is isomorphic to $\nu_{M}$ (as an unoriented bundle), which implies (2). ∎
Theorem 1.5 follows immediately from this lemma and Theorem 1.1.
Example 4.2 (Dehn surgeries on the trefoil knot).
Let $K=T(2,3)$ be the right-handed trefoil knot, and let $\mu$ and $\lambda$ denote the standard meridian and longitude of $K$. There is a natural Seifert fibre structure on the exterior $X=S^{3}\setminus\nu(K)$, whose regular fibre has slope $6\mu+\lambda$ on $\partial X$. Hence, $p/q$-surgery on $K$ yields a small Seifert fibred space whenever $p/q\neq 6$.
It is known that $X(p/q)$ admits co-oriented taut foliations when $p/q<1$ [16, §3]. On the other hand, [9, Theorem 1.7] shows that the Euler class of the co-oriented taut foliation constructed by Roberts in [16] vanishes if and only if
$|q|\equiv 1\pmod{p}$.
Combining Example 4.2 with Lemma 4.1 yields Theorem 1.6.
The following lemma is an immediate consequence of Theorem 1.1.
Lemma 4.3.
If $e(\nu_{M})=0$, then $d=\gcd(b_{i},b_{j})\implies a_{i}\equiv a_{j}\mbox{ {\rm(mod $d$)}}$.
Two sets of Seifert invariants $(b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$, $(b^{\prime},\frac{a_{1}^{\prime}}{b_{1}},\ldots,\frac{a_{n}^{\prime}}{b_{n}})$ are said to be equivalent, written
$(b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})\sim(b^{\prime},\frac{a_{1}^{\prime}}{b_{1}},\ldots,\frac{a_{n}^{\prime}}{b_{n}})$
if they differ by a finite sequence of moves each of which either multiplies the coordinates by 1-1 or is of the form described in Remark 2.1(1). Equivalent sets of Seifert invariants yield homeomorphic manifolds.
4.1. $\mathcal{B}$ is spherical
When $\mathcal{B}$ is spherical, $n=3$ and $(b_{1},b_{2},b_{3})$ is one of the Platonic triples $(2,2,b_{3})$, $(2,3,3)$, $(2,3,4)$, or $(2,3,5)$. In each case, $b_{1}=2$ and $b_{2}$ is either $2$ or $3$. The fundamental group of $M$ is finite in all cases and therefore $b_{1}b_{2}b_{3}e(M)=\pm|H_{1}(M)|\neq 0$ (i.e., $H_{1}(M)$ is finite). Further, $\chi(\mathcal{B})>0$.
By assumption, $a_{i}=1$ when $b_{i}=2$. If $b_{2}=3$ and $a_{2}=2$, then the Seifert invariants of $M$ are
$(b,\frac{1}{2},\frac{2}{3},\frac{a_{3}}{b_{3}})\sim(b+1,\frac{1}{2},\frac{-1}{3},\frac{a_{3}}{b_{3}})\sim(-b-1,-\frac{1}{2},\frac{1}{3},-\frac{a_{3}}{b_{3}})\sim(-b-3,\frac{1}{2},\frac{1}{3},\frac{b_{3}-a_{3}}{b_{3}})$
Thus we can assume that $M$ has Seifert invariants $(b,\frac{1}{2},\frac{1}{b_{2}},\frac{a}{b_{3}})$. The following table lists $\chi(\mathcal{B}),e(M)$ and $|H_{1}(M)|$ for these manifolds. The last column determines the Seifert invariants of $M$ when $e(\nu_{M})=0$ by applying Theorem 1.1 and Corollary 1.3.
Seifert invariants $\chi(\mathcal{B})$ $e(M)$ $\pm|H_{1}(M)|$ Seifert invariants when $e(\nu_{M})=0$
$(b,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$ $\tfrac{1}{2}$ $\tfrac{2b+3}{2}$ $4(2b+3)$ $(-1,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$ and $(-2,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$
$(b,\tfrac{1}{2},\tfrac{1}{2},\tfrac{a}{b_{3}}),\,b_{3}\geq 3$ $\tfrac{1}{b_{3}}$ $\tfrac{b_{3}(b+1)+a}{b_{3}}$ $4(b_{3}(b+1)+a)$ $(-1,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{b_{3}})$
$(b,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{3})$ $\tfrac{1}{6}$ $\tfrac{6b+7}{6}$ $3(6b+7)$ $(-1,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{3})$
$(b,\tfrac{1}{2},\tfrac{1}{3},\tfrac{a}{4})$ $\tfrac{1}{12}$ $\tfrac{12b+10+3a}{12}$ $2(12b+10+3a)$ $(-1,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{4})$
$(b,\tfrac{1}{2},\tfrac{1}{3},\tfrac{a}{5})$ $\tfrac{1}{30}$ $\tfrac{30b+6a+25}{30}$ $30b+6a+25$ $(-1,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{5})$
It is interesting to note that the values of the Seifert invariants when $e(\nu_{M})=0$ correspond to those which realise the minimal values of $|H_{1}(M)|$ in the given families.
4.2. $\mathcal{B}$ is Euclidean
Here, either $n=3$ and $(b_{1},b_{2},b_{3})$ is one of the triples $(2,3,6),(2,4,4)$, $(3,3,3)$ or $n=4$ and $(b_{1},b_{2},b_{3},b_{4})=(2,2,2,2)$. We can assume that $a_{i}=1$ in all cases. This is obvious when $b_{i}=2$ and using Lemma 4.3 and the arguments of §4.1 deals with the case that some $b_{i}>2$.
As $\mathcal{B}$ is Euclidean, its Euler characteristic is zero. Hence if $\nu_{M}$ is trivial, Corollary 1.3 implies that $e(M)=0$ and we can use this to determine the Seifert invariants of $M$ when $e(\nu_{M})=0$.
Seifert invariants $e(M)$ $\pm|H_{1}(M)|$ Seifert invariants when $e(\nu_{M})=0$
$(b,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{6})$ $b+1$ $36(b+1)$ $(-1,\tfrac{1}{2},\tfrac{1}{3},\tfrac{1}{6})$
$(b,\tfrac{1}{2},\tfrac{1}{4},\tfrac{1}{4})$ $b+1$ $32(b+1)$ $(-1,\tfrac{1}{2},\tfrac{1}{4},\tfrac{1}{4})$
$(b,\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$ $b+1$ $27(b+1)$ $(-1,\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$
$(b,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$ $b+2$ $16(b+2)$ $(-2,\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$
As in the spherical case, those $M$ with $e(\nu_{M})=0$ are characterised by their first homology groups having minimal order (0 in this case). Their first Betti numbers are positive and therefore they admit horizontal fibrations (by tori) over the circle. This proves Theorem 1.7 when $\mathcal{B}$ is Euclidean.
4.3. $\mathcal{B}$ is hyperbolic
We assume in this section that the Seifert invariants of
$M=M(0;b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$
have been normalised so that $0<a_{i}<b_{i}$ for $1\leq i\leq n$.
The results of [12] and [15] show that $M$ admits a horizontal foliation if and only if one of the following conditions holds:
  1. (1)
    $2-n\leq b\leq-2$;
  2. (2)
    b=1b=-1 and there are coprime integers $0<c<d$ such that for some permutation $(\frac{c_{1}}{d},\frac{c_{2}}{d},\ldots,\frac{c_{n}}{d})$ of $(\frac{c}{d},\frac{d-c}{d},\frac{1}{d},\ldots,\frac{1}{d})$ we have $\frac{a_{i}}{b_{i}}<\frac{c_{i}}{d}$ for each $i$.
  3. (3)
    b=1nb=1-n and there are coprime integers $0<c<d$ such that for some permutation $(\frac{c_{1}}{d},\frac{c_{2}}{d},\ldots,\frac{c_{n}}{d})$ of $(\frac{c}{d},\frac{d-c}{d},\frac{1}{d},\ldots,\frac{1}{d})$ we have $\frac{b_{i}-a_{i}}{b_{i}}<\frac{c_{i}}{d}$ for each $i$.
The following specific examples illustrate that there are many Seifert manifolds $M$ that have co-orientable taut foliations with non-zero Euler class.
Examples 4.4.
Using Lemma 4.3, if some $b_{i},b_{j}$ have a common factor greater than $2$, one can construct many examples $M$ with $e(\nu_{M})\neq 0$. On the other hand using the conditions (1), (2), (3) immediately above, one can arrange for many of these to have horizontal foliations.
In (1), (2), (3), and (4) below we give some other classes of examples of Seifert fibred manifolds which admit co-orientable taut foliations but none with vanishing Euler class. The proofs are all applications of Theorem 1.1; we omit the details.
(1) By Lemma 4.3, if $b_{1}=b_{2}=\cdots=b_{n}$ and $e(\nu_{M})=0$, then $a_{1}=a_{2}=\cdots=a_{n}$. The first set of examples deals with a special case of this situation.
For $n\geq 4$, consider the manifold $M=M(0;b,\frac{1}{c},...,\frac{1}{c})$. Then
  • (a)
    $M$ has a horizontal foliation if and only if either $2-n\leq b\leq-2$, or b=1b=-1 and $c\geq 3$.
  • (b)
    If $c\geq 2n-2$, then $e(\nu_{M})=0$ if and only if b=2nb=2-n.
The bound on $c$ in part (b) is optimal: if c=2n3c=2n-3 and b=1b=-1, then $e(\nu_{M})=0$.
(2) For $n\geq 3$, let $M=M(0;2-n,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$, where $\sum_{i}\frac{a_{i}}{b_{i}}\leq\frac{n-2}{2}$. Then $M$ has a horizontal foliation, but $e(\nu_{M})=0$ if and only if $a_{i}=1$ for $1\leq i\leq n$.
(3) For $n\geq 4$, consider the manifold $M=M(0;b,\frac{1}{b_{1}},\ldots,\frac{1}{b_{n}})$, where the $b_{i}$ are pairwise coprime and $\geq n$. Then
  • (a)
    $M$ has a horizontal foliation if and only if $2-n\leq b\leq-1$.
  • (b)
    $e(\nu_{M})=0$ if and only if b=2nb=2-n.
(4) For $n\geq 4$, let $M=M(0;b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$, where $\sum_{i}\frac{a_{i}}{b_{i}}\leq\frac{n}{6}$. Then
  • (a)
    $M$ has a horizontal foliation for $2-n\leq b\leq-2$, and for b=1b=-1 if $n=4,5$, or $6$.
  • (b)
    $e(\nu_{M})=0$ for at most one $b$ with $2-n\leq b\leq\frac{2-n}{2}$.
4.4. Small values of $|H_{1}(M)|$
It is known that if $|H_{1}(M)|\leq 2$, then the Euler class $e(\nu_{M})$ vanishes (see [9, Proposition 2.1] for instance). In this section, we show that this fact can also be obtained from the computations in this article.
We have already discussed the case $M=M(0;b)$ in Remark 2.2, so we assume that $n>0$.
We recall the presentation of $H^{2}(\pi_{1}(\mathcal{B}))$ in §3.1, where $H^{2}(\pi_{1}(\mathcal{B}))$ is generated by $A_{0},A_{1},\ldots,A_{n}$ with relations $A_{0}=b_{i}A_{i}$ for $i=1,\ldots,n$. In §3, we showed that $e(\nu_{M})=\alpha^{*}(e(\overline{\varphi}))=0$, where $\overline{\varphi}\colon\pi_{1}(\mathcal{B})\to PSL(2,\mathbb{R})$ is discrete and faithful, if and only if $\exists\,m\in\mathbb{Z}$ such that $e(\overline{\varphi})=m\,e(E_{M})$. See (3.1.2) and (3.1.3). When $M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, by (3.1.1) and (3.2.2),
$e(E_{M})=-\bigl(bA_{0}+a_{1}A_{1}+\cdots+a_{n}A_{n}\bigr),$
and
$e(\overline{\varphi})=-\left((2-n)A_{0}+A_{1}+A_{2}+\cdots+A_{n}\right)$
Suppose that $m$ is an integer such that $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i})$ for $i=1,\cdots,n$. Set
$S_{0}:=m(bA_{0}+a_{1}A_{1}+\cdots+a_{n}A_{n})=(mb+\sum_{i=1}^{n}\left\lfloor\frac{ma_{i}}{b_{i}}\right\rfloor)A_{0}+A_{1}+\cdots+A_{n}$
Any other integer $m^{\prime}$ for which the congruences $m^{\prime}a_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n$ hold can be written $m^{\prime}=m+rl$, where $l$ is the least common multiple of $b_{1},b_{2},\ldots,b_{n}$ and $r\in\mathbb{Z}$. Set $m_{r}=m+rl$. The reader will verify that
(4.4.1) $S_{r}:=m_{r}(bA_{0}+a_{1}A_{1}+\cdots+a_{n}A_{n})=S_{0}+rle(M)A_{0}$
We noted after the statement of Corollary 1.3 that $e(M)=\varepsilon|H_{1}(M)|/b_{1}b_{2}\cdots b_{n}$ for some $\varepsilon\in\{\pm 1\}$. Thus from (4.4.1) we have
(4.4.2) $S_{r}=S_{0}+\varepsilon r\frac{|H_{1}(M)|}{(b_{1}b_{2}\cdots b_{n}/l)}A_{0}$
When $M$ is an integer homology $3$-sphere, $H^{2}(M)=0$, so $\nu_{M}$ is trivial. This is reflected in the fact that $l=b_{1}b_{2}\cdots b_{n}$ in this case, so we can choose $r$ so that the coefficient of $A_{0}$ in $S_{r}$ is 2n2-n.
Similarly, if $H_{1}(M)\cong\mathbb{Z}/2$, the fact that the tangent bundle to the Seifert fibres of $M$ is a trivial line bundle combines with the method of proof of [9, Proposition 2.1] to show that $e(\nu_{M})=0$. We leave it to the reader to verify this using (4.4.2) and Theorem 1.1.

### v2

@@
-The Euler class of the normal bundle of a Seifert fibration and horizontal foliations
+On $3$-manifolds admitting co-orientable taut foliations, but none with vanishing Euler class
 Abstract.
-For Seifert fibred manifolds with orientable base orbifolds, we establish a necessary and sufficient condition for the Euler class of the normal bundle of the Seifert fibration to vanish.
-When the base orbifold is hyperbolic, we also provide a second proof of this condition from the perspective of discrete faithful representations of Fuchsian groups.
-As an application, we present infinitely many Seifert fibred rational homology spheres that admit co-oriented taut foliations but none with vanishing Euler class.
-In the context of the $L$-space conjecture, these provide examples of rational homology spheres that admit co-oriented taut foliations (and hence are not $L$-spaces) and have left-orderable fundamental groups yet none of the left-orders arise directly from the universal circle actions associated to co-oriented taut foliations.
+In this article, we construct infinitely many (small Seifert fibred, hyperbolic and toroidal) rational homology $3$-spheres that admit co-orientable taut foliations, but none with vanishing Euler class.
+In the context of the $L$-space conjecture, these examples provide rational homology $3$-spheres that admit co-orientable taut foliations (and hence are not $L$-spaces) and have left-orderable fundamental groups, yet none of the left orders arise directly from the universal circle actions associated to co-orientable taut foliations.
+The hyperbolic and non-Seifert toroidal examples are obtained from Dehn surgeries on knots in the $3$-sphere and use Heegaard Floer homology to obstruct the existence of a co-orientable foliation with vanishing Euler class.
+For the Seifert fibred case, we establish necessary and sufficient conditions for the Euler class of the normal bundle of the Seifert fibration to vanish.
+Moreover, when the base orbifold is hyperbolic, we also provide a second proof of this condition from the viewpoint of discrete faithful representations of Fuchsian groups.
 Steven Boyer was partially supported by NSERC grant RGPIN 2024-06196
 Ying Hu was partially supported by NSF grant DMS-2409398
+Duncan McCoy was partially supported by NSERC grant RGPIN-2020-05491 and a Canada Research Chair
 2010 Mathematics Subject Classification. Primary 57M25, 57M50, 57M99
-Key words: Seifert fibred manifold, normal bundle, Euler class, horizontal foliation
+Key words: taut foliation, Knot Floer complex, self-conjugate spinc structure, Seifert fibred manifold, normal bundle, Euler class, horizontal foliation.
 1. Introduction
-In this article, we study the triviality of the normal bundle $\nu_{M}$ to the Seifert fibres of a Seifert fibration on a $3$-manifold $M$ with orientable base orbifold.
-We will refer to $\nu_{M}$ as the normal bundle of the Seifert fibration.
-As an oriented $2$-plane bundle, $\nu_{M}$ is trivial if and only if its Euler class $e(\nu_{M})\in H^{2}(M)$ vanishes.
-We denote a closed oriented Seifert fibred manifold $M$ with orientable base orbifold by $M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, where $b\in\mathbb{Z}$ and $b_{i}\geq 2$ for $i=1,\ldots,n$.
+1.1. Rational homology spheres admitting no taut foliations with vanishing Euler class
+It is predicted by the $L$-space conjecture [4, 18] that if a closed, orientable, prime $3$-manifold $M$ admits a co-orientable taut foliation, then its fundamental group is left-orderable.
+One of the main tools for proving this is to use Thurston’s universal circle action, which to each co-oriented taut foliation $\mathcal{F}$ on an oriented rational homology $3$-sphere $M$ associates a non-trivial representation
+$\rho_{\mathcal{F}}:\pi_{1}(M)\to\mbox{Homeo}_{+}(S^{1}).$
+See [10] for the details.
+If one can show that this representation lifts to the universal covering group $\mbox{Homeo}_{\mathbb{Z}}(\mathbb{R})$ of $\mbox{Homeo}_{+}(S^{1})$, then it follows from [8, Theorem 1.1] that the fundamental group of $M$ is left-orderable.
+However, such a lift exists if and only if the Euler class of $\rho_{\mathcal{F}}$ vanishes.
+Since this Euler class coincides with the Euler class of the tangent bundle of $\mathcal{F}$ [5], which we denote by $e(\mathcal{F})$, we are naturally led to the following question.
+Question 1.1.
+If a closed, orientable, prime $3$-manifold admits a co-orientable taut foliation, does it admit one whose Euler class is zero?
+[^1] Given a co-orientable taut foliation, the Euler class of its tangent bundle is well-defined up to sign.
+So it makes sense to discuss if a co-orientable taut foliation has vanishing Euler class.
+Although it has been expected that the answer to the question is no, we do not know of any examples verifying this. In this article, we show that such examples occur quite frequently.
+Convention: When referring to $p/q$-surgery on a knot, we assume that $p$ and $q$ are coprime integers with $p\geq 0$.
+Theorem 1.2.
+Let $M$ be obtained by $p/q$-surgery on a knot $K$ in the $3$-sphere. If $q$ is even and |p/q|>2g(K)1|p/q|>2g(K)-1, then $M$ does not admit a co-orientable taut foliation with Euler class zero.
+Thus the existence of a knot for which some large surgery with even denominator admits a co-orientable taut foliation shows that the answer to Question 1.1 is no.
+For instance, Roberts has shown that if $K$ is a non-trivial fibred knot in the $3$-sphere, then $K(p/q)$, the manifold obtained by $p/q$-surgery on $K$, admits a co-orientable taut foliation for all $p/q$, $p/q<1$, or p/q>1p/q>-1 depending on whether the fractional Dehn twist coefficient of $K$ is zero, positive, or negative respectively [26].
+The following corollary states the case when the fractional Dehn twist coefficient is zero.
+Examples of such knots include all hyperbolic fibred alternating knots [6, 2].
+Corollary 1.3.
+Suppose that $K$ is a non-trivial fibred knot in the $3$-sphere whose fractional Dehn twist coefficient is zero.
+Then for all slopes $p/q$ with $q$ even and |p/q|>2g(K)1|p/q|>2g(K)-1, $K(p/q)$ admits a co-orientable taut foliation, but none with zero Euler class.
+A similar conclusion holds for many non-fibred knots.
+Indeed, the $L$-space conjecture predicts that every irreducible rational surgery on a non-fibred knot admits a co-orientable taut foliation and this is known to be true in many instances.
+For instance, Roberts and Delman proved this for non-torus alternating knots and for Montesinos knots which are not $L$-space knots.
+Examples of manifolds having left-orderable fundamental groups and admitting co-orientable taut foliations, but none with vanishing Euler class, can also be constructed, thus showing that no left-orders on the group arise from lifts of universal circle representations.
+For instance, the figure eight knot is a fibred hyperbolic knot of genus one with zero fractional Dehn twist coefficient and each rational surgery on it has a left-orderable fundamental group [27].
+Corollary 1.4.
+Suppose that $K$ is the figure eight knot.
+Then for all slopes $p/q$ with $q$ even and $|p/q|>1$, $K(p/q)$ has a left-orderable fundamental group and admits a co-orientable taut foliation, but no co-orientable taut foliation with zero Euler class.
+By way of contrast, we note that Theorem 1.2 cannot be extended to all large surgeries on a knot without some hypotheses on the slope $p/q$.
+For instance, it is known that $p$-surgery on the figure eight knot admits co-orientable taut foliations of zero Euler class for each integer $p$ [11, 14].
+More generally, we have
+Proposition 2.3.
+Let $M$ be a rational homology sphere obtained by integer surgery on a genus one knot in the $3$-sphere. Then every co-orientable taut foliation on $M$ has Euler class zero.
+Remark 1.5.
+The line of reasoning used in the proof of Proposition 2.3 can easily be extended to show that if $K$ is an almost $L$-space knot, meaning a knot with positive almost $L$-space surgeries [9], and $M$ is obtained by $p$-surgery on $K$ for some $p\geq 2g(K)-1$, then any co-orientable taut foliation on $M$ has Euler class zero.
+This shows that for knots of arbitrarily large genus, constraints must be placed on the denominators of surgery slopes for Theorem 1.2 to hold.
+Corollary 1.6.
+Suppose that $K$ is a knot of genus one in the $3$-sphere.
+If $p$ is an integer such that $K(p)$ admits a co-orientable taut foliation, then $\pi_{1}(K(p))$ is left-orderable.
+Similarly, if $K$ is an almost $L$-space knot and $p\geq 2g(K)-1$ is an integer such that $K(p)$ admits a co-orientable taut foliation, then $\pi_{1}(K(p))$ is left-orderable.
+All the examples in Corollary 1.4 are hyperbolic.
+The simplest examples of Seifert fibred manifolds in which every co-orientable taut foliation has non-zero Euler class arise from Dehn surgery on the trefoil knot.
+Theorem 1.7.
+Suppose that $K$ is the right-handed trefoil knot.
+Then $K(p/q)$ admits a co-orientable taut foliation and has a left-orderable fundamental group if and only if $p/q<1$.
+It admits a co-orientable taut foliation with zero Euler class if and only if $p/q<1$ and $|q|\equiv 1\pmod{p}$.
+In general, we denote a closed oriented Seifert fibred manifold $M$ with orientable base orbifold by $M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, where $b\in\mathbb{Z}$ and $b_{i}\geq 2$ for $i=1,\ldots,n$.
 See §3.1 for a detailed explanation of the notation.
@@
 Recall that the Euler characteristic $\chi(\mathcal{B})$ of $\mathcal{B}$ is $\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$, and the Euler number of $M$ is $e(M)=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}$.
-Here is our main result.
+Theorem 1.8.
+Suppose that $M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ is a rational homology $3$-sphere which admits a co-orientable taut foliation $\mathcal{F}$.
+Then $\mathcal{F}$ has zero Euler class if and only if there exists $m\in\mathbb{Z}$ such that
+  1. (1)
+    $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n,\;\text{and}$
+  2. (2)
+    $me(M)=\chi(\mathcal{B})$.
+We can also produce toroidal examples which are not Seifert fibred.
+Define the JSJ graph of a knot $K$ to be the rooted tree dual to the JSJ tori of the exterior $X(K)$ of $K$, where the root vertex corresponds to the piece of the JSJ decomposition containing $\partial X(K)$.
+Any finite rooted tree can arise as the JSJ graph of a knot.
+We say that a JSJ graph is a rooted interval when it is homeomorphic to an interval with root at an endpoint.
+Since non-integral surgeries on knots in $S^{3}$ are irreducible [12], the following corollary follows immediately from Theorem 1.2 and Theorems 3.13 and 4.3 of [3].
+Corollary 1.9.
+Suppose that $K$ is a fibred satellite knot whose $JSJ$ graph is not a rooted interval.
+Then for all slopes $p/q$ with $q$ even and |p/q|>2g(K)1|p/q|>2g(K)-1, $K(p/q)$ is toroidal, has a left-orderable fundamental group, and admits a co-orientable taut foliation, but none with zero Euler class.
+Remark 1.10.
+Fibred composite knots satisfy the corollary’s hypothesis. Taking $K$ to be a connected sum of torus knots yields examples which are graph manifolds.
+1.2. Euler class of the normal bundle of a Seifert fibration
+While Theorem 1.2 is proved using the knot Floer complex, the key fact in analyzing the Euler classes of co-oriented taut foliations on a Seifert fibred rational homology $3$-sphere $M$ is that, up to isotopy, all such foliations are horizontal; that is, transverse to the Seifert fibres of $M$ (Lemma 5.1).
+Hence, the tangent bundle of $\mathcal{F}$ is isomorphic to the normal bundle $\nu_{M}$ to the Seifert fibres of $M$, while the latter makes sense even without the existence of a horizontal foliation.
+We will refer to $\nu_{M}$ as the normal bundle of the Seifert fibration. In §3 and §4, we determine the following necessary and sufficient conditions for its Euler class to vanish.
 Theorem 1.11.
@@
     $me(M)=\chi(\mathcal{B})$.
-Remark 1.2.
-It is often easy to identify an initial $m\in\mathbb{Z}$ that satisfies the congruence equations $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i})$ for $i=1,\cdots,n$.
-Any other integer $m^{\prime}$ satisfying these congruences can be written as $m^{\prime}=m+rl$, where $l$ is the least common multiple of $b_{1},b_{2},\ldots,b_{n}$ and $r\in\mathbb{Z}$.
-Hence when $n>0$, $e(M)\neq 0$ and $\chi(\mathcal{B})\neq 0$, the condition $m\,e(M)=\chi(\mathcal{B})$ in Theorem 1.1 can be replaced by
-$\frac{\chi(\mathcal{B})}{e(M)}\text{ is an integer congruent to }m\;(\operatorname{mod}l)$
+Theorem 1.8 follows from Theorem 1.11 and Lemma 5.1.
 Next, we deduce the following easily verifiable necessary condition for $\nu_{M}$ to be a trivial bundle:
 Corollary 1.12.
-Given a closed oriented Seifert fibred manifold $M$, if the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes, then either $M$ is an oriented $S^{1}$-bundle over the torus or one of the following holds:
+Given a closed oriented Seifert fibred manifold $M$, if the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes, then either $M$ is an orientable $S^{1}$-bundle over the torus or one of the following holds:
   1. (1)
@@
     $\frac{\chi(\mathcal{B})}{e(M)}\in\mathbb{Z}.$
-Note that if $M$ is an oriented $S^{1}$-bundle over the torus, then its normal bundle is always trivial (see Remark 2.2).
+Note that if $M$ is an orientable $S^{1}$-bundle over the torus, then its normal bundle is always trivial (see Remark 3.2).
 By abelianizing the presentation (4.0.1) of $\pi_{1}(M)$ obtained by viewing $M$ as doing a Dehn filling of a trivial $S^{1}$-bundle over an $n$-punctured orientable surface of genus $g$ as in §3.1, one can show that if $T_{1}(M)$ denotes the torsion subgroup of $H_{1}(M)$, then
@@
 Hence, if $M$ is a rational homology $3$-sphere, then $b_{1}\ldots b_{n}\chi(\mathcal{B})$ is a multiple of $|H_{1}(M)|$.
-Our initial motivation for studying the Euler class of $\nu_{M}$ comes from the study of horizontal foliations on Seifert fibred manifolds.
-By definition, a horizontal foliation is transverse to the Seifert fibration and as such, their tangent bundles are isomorphic to $\nu_{M}$.
-Hence, their Euler classes coincide, at least up to sign.
-It is predicted by the $L$-space conjecture [3, 13] that if a closed, orientable, prime $3$-manifold admits a co-orientable taut foliation, then its fundamental group is left-orderable.
-One of the main tools for proving left-orderability of the fundamental group in the presence of a co-orientable taut foliation is to use Thurston’s universal circle action.
-See [7, 4] for the details.
-However, this approach requires the Euler class of the taut foliation to vanish.
-Although it was expected that examples of $3$-manifolds admitting co-orientable taut foliations but none with vanishing Euler class should exist, no published examples were previously known.
-In this article, we settle the matter by presenting such examples within the family of Seifert fibred rational homology $3$-spheres.
-Our arguments are based on the following application of Theorem 1.1.
-Theorem 1.5.
-Suppose that $M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ is a rational homology $3$-sphere which admits a co-orientable taut foliation $\mathcal{F}$.
-Then $\mathcal{F}$ has zero Euler class if and only if there exists $m\in\mathbb{Z}$ such that
-  1. (1)
-    $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n,\;\text{and}$
-  2. (2)
-    $me(M)=\chi(\mathcal{B})$.
-The simplest examples of manifolds in which every co-oriented taut foliation has non-zero Euler class arise from Dehn surgery on the trefoil knot.
-In the following statement, we assume that $p$ and $q$ are nonzero coprime integers.
-Theorem 1.6.
-The $3$-manifold obtained by $p/q$-surgery along the right-handed trefoil knot admits a co-oriented taut foliation if and only if $p/q<1$.
-It admits one with zero Euler class if and only if $p/q<1$ and $|q|\equiv 1\pmod{p}$.
-We will present more examples of rational homology Seifert fibred manifolds that admit co-oriented taut foliations, but none with vanishing Euler class, in §4.3 (Examples 4.4).
-In fact, our computations in §4 suggests that most closed Seifert fibred manifolds have nontrivial normal bundle, including the case when the base orbifold is Euclidean (i.e.
+We will present more examples of rational homology Seifert fibred manifolds that admit co-orientable taut foliations, but none with vanishing Euler class, in §5.3 (Examples 5.3).
+In fact, our computations in §5 suggest that most closed Seifert fibred manifolds have nontrivial normal bundle, including the case when the base orbifold is Euclidean (i.e.
 $\chi(\mathcal{B})=0$).
@@
 Plan of the paper
-In Section 2 we explain our notation for Seifert manifolds and prove Theorem 1.1, which gives necessary and sufficient conditions for the Euler class of the normal bundle of a Seifert fibration to vanish.
+We prove Theorem 1.2 and Proposition 2.3 in §2.
+In §3 we explain our notation for Seifert manifolds and prove Theorem 1.11, which gives necessary and sufficient conditions for the Euler class of the normal bundle of a Seifert fibration to vanish.
 We also prove Corollary 1.12 in this section.
-In Section 3 we present a different perspective on Theorem 1.1 and Corollary 1.3 in the case that the base orbifold is hyperbolic through the use of discrete faithful $PSL(2,\mathbb{R})$-representations of $\pi_{1}(\mathcal{B})$.
+In §4 we present a different perspective on Theorem 1.11 and Corollary 1.12 in the case that the base orbifold is hyperbolic through the use of discrete faithful $PSL(2,\mathbb{R})$-representations of $\pi_{1}(\mathcal{B})$.
 As a consequence of this approach, we prove the hyperbolic case of Theorem 1.14.
-Finally, in Section 4 we prove Theorem 1.5 and Theorem 1.6 and provide examples of Seifert fibred manifolds whose normal bundles are (non-)trivial.
+Finally, in §5 we prove Theorems 1.7 and 1.8, and provide examples of Seifert fibred manifolds whose normal bundles are (non-)trivial.
 We classify which occurs when $\mathcal{B}$ is spherical in §5.1 and is Euclidean in §5.2.
@@
 The case that $\mathcal{B}$ is hyperbolic is analysed in §5.3.
-Finally, in §4.4, we illustrate how the conditions of Theorem 1.1 can be used to verify the known fact that $e(\nu_{M})=0$ when $|H_{1}(M)|=1,2$.
 Acknowledgements
@@
 The second and third named authors would like to thank the Simons Foundation, the CRM, and CIRGET for their support during the preparation of this article.
+2. Examples from surgeries on knots in the $3$-sphere
+Theorem 1.2 is an immediate consequence of the following two lemmas. The first follows from Ozsváth and Szabó’s proof that a manifold with a co-orientable taut foliation is not an $L$-space [22].
+Lemma 2.1.
+Let $M$ be a rational homology $3$-sphere admitting a co-orientable taut foliation $\mathcal{F}$.
+Then $M$ has a $\mathrm{spin}^{c}$-structure $\mathfrak{s}$ such that $HF_{\mathrm{red}}(M,\mathfrak{s})\neq 0$ and $c_{1}(\mathfrak{s})=e(\mathcal{F})$.
+Proof.
+Let $M$ be a rational homology $3$-sphere that admits a co-orientable taut foliation $\mathcal{F}$.
+The tangent plane field of $\mathcal{F}$ can be perturbed to obtain a tight contact structure $\xi$ whose contact invariant $c(\xi)\neq 0\in HF_{\mathrm{red}}(M,\mathfrak{s})$, with $c_{1}(\mathfrak{s})=e(\mathcal{F})$ [22].
+∎
+The second is an application of the mapping cone formula for calculating the Heegaard Floer homology of surgeries on a knot in the $3$-sphere [23, 24].
+In conjunction with Lemma 2.1, this implies Theorem 1.2 since for a 3-manifold $M$ for which $H_{1}(M)$ has odd order, the unique spinc-structure satisfying $c_{1}(\mathfrak{s})=0$ is also the unique self-conjugate spinc-structure on $M$.
+Lemma 2.2.
+Let $M$ be obtained by $p/q$-surgery on a knot $K$ in the $3$-sphere, where $q$ is even and |p/q|>2g(K)1|p/q|>2g(K)-1.
+Then the unique self-conjugate $\mathrm{spin}^{c}$-structure $\mathfrak{s}$ on $M$ satisfies $HF_{\mathrm{red}}(M,\mathfrak{s})=0$.
+Proof.
+Without loss of generality we may assume that $p/q>0$.
+Let $CFK^{\infty}(K)=\bigoplus_{i,j\in\mathbb{Z}}C\{(i,j)\}$ be the knot Floer complex for $K$.
+Since |p/q|>2g(K)1|p/q|>2g(K)-1, the mapping cone formula for computing $HF^{+}(M)$ in terms of $CFK^{\infty}(K)$ implies that there is a labelling
+$\phi:\mathbb{Z}/p\to\mathrm{Spin}^{c}(M),$
+such that [21, Proposition 3.5]
+$HF_{\mathrm{red}}(M,\phi(i))\cong\bigoplus_{k\in\mathbb{Z}}\bm{A}_{{\rm red},\,\lfloor\frac{pk+i}{q}\rfloor}(K),$
+where $\bm{A}_{{\rm red},\,k}(K)$ denotes the reduced part of
+$\bm{A}^{+}_{k}=H_{*}(C\{i\geq 0\text{ or }j\geq k\}).$
+However, $\bm{A}_{{\rm red},\,k}(K)=0$ for |k|>g(K)1|k|>g(K)-1 [21, Lemma 3.2]. Thus, for $0\leq i\leq p-1$, this becomes
+$HF_{\mathrm{red}}(M,\phi(i))\cong\begin{cases}\bm{A}_{{\rm red},\,\lfloor\frac{i}{q}\rfloor}(K),&0\leq i<qg(K)-q,\\ \bm{A}_{{\rm red},\,\lfloor\frac{i-p}{q}\rfloor}(K),&p-qg(K)<i<p,\\ 0,&qg(K)-q\leq i\leq p-qg(K).\end{cases}$
+In particular, for each $|j|\leq g(K)-1$, there are precisely $q$ $\mathrm{spin}^{c}$-structures for which $HF_{\mathrm{red}}(M,\mathfrak{s})$ is isomorphic to $\bm{A}_{{\rm red},\,j}(K)$, and for all other $\mathrm{spin}^{c}$-structures, $HF_{\mathrm{red}}(M,\mathfrak{s})=0$.
+Thus the set
+$S=\{\mathfrak{s}\in\mathrm{Spin}^{c}(M)\mid HF_{\mathrm{red}}(M,\mathfrak{s})\neq 0\}$
+has cardinality divisible by $q$.
+In particular, $|S|$ is even.
+Since conjugation of $\mathrm{spin}^{c}$-structures induces an isomorphism of Heegaard Floer homology, conjugation acts as an involution on $S$.
+This implies that there are an even number of self-conjugate $\mathrm{spin}^{c}$-structures in $S$.
+Since $p$ is odd, there is precisely one self-conjugate spinc-structure on $M$, so it follows that $S$ does not contain the self-conjugate spinc-structure on MM.
+[^2] An alternative line of argument is to use the fact that conjugation of spinc-structures corresponds to the involution $i\mapsto p+q-1-i$ (mod $p$) on $\mathbb{Z}/p$.
+For $q$ even, this implies that the self-conjugate spinc-structure is labelled by $s=\frac{p+q-1}{2}$.
+However since |p/q|>2g(K)1|p/q|>2g(K)-1, the above calculation shows that $HF_{\mathrm{red}}(M,\phi(s))=0$.
+In the opposite direction, we note that Theorem 1.2 cannot be extended to all large surgeries on a knot $K$ without some hypotheses on the slope $p/q$.
+Proposition 2.3.
+Let $M$ be a rational homology sphere obtained by integer surgery on a genus one knot in the 3-sphere. Then every co-orientable taut foliation on $M$ has Euler class zero.
+Proof.
+Let $K$ be a genus one knot and let $M$ be obtained by $p$-surgery on $K$.
+Without loss of generality we may assume that $p>0$.
+The proposition is deduced from Lemma 2.1 by showing that there is at most one spinc-structure $\mathfrak{s}_{0}$ for which $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\neq 0$ and then verifying that $c_{1}(\mathfrak{s}_{0})=0$.
+First we compute $HF_{\mathrm{red}}(M)$. For integer surgeries of slope $p\geq 2g(K)-1$, the mapping cone formula gives a labelling
+(2.0.1) $\phi:\mathbb{Z}/p\to\mathrm{Spin}^{c}(M),$
+such that [21, Proposition 3.5]
+$HF_{\mathrm{red}}(M,\phi(i))\cong\bigoplus_{k\in\mathbb{Z}}\bm{A}_{{\rm red},\,i+kp}(K).$
+However, as $g(K)=1$, we have $\bm{A}_{{\rm red},\,k}(K)=0$ for $|k|>0$ [21, Lemma 3.2]. Thus, for $0\leq i\leq p-1$, this becomes
+$HF_{\mathrm{red}}(M,\phi(i))\cong\begin{cases}\bm{A}_{{\rm red},\,0}(K),&i=0,\\ 0,&0<i\leq p-1,\end{cases}$
+giving that that there is at most one spinc-structure $\mathfrak{s}_{0}$ for which $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\neq 0$ and that this is the spinc-structure corresponding to 0 under (2.0.1).
+We verify that $c_{1}(\mathfrak{s}_{0})=0$ by unpacking the labelling in (2.0.1).
+As in [23, Section 2.4] the map $\phi$ is obtained as follows.
+Take $X$ to be the trace of $p$-surgery on $K$.
+Let $\Sigma]\in H_{2}(X)$ denote that the homology class obtained by capping off the core of the 2-handle with a Seifert surface for $K$.
+If $\mathfrak{s}\in\mathrm{Spin}^{c}(M)$ is the restriction of $\mathfrak{t}\in\mathrm{Spin}^{c}(X)$, then $\phi(i)=\mathfrak{s}$, where $i$ satisfies
+$\langle c_{1}(\mathfrak{t}),[\Sigma]\rangle+p\equiv 2i\bmod 2p.$
+Thus the spinc-structure with $\mathfrak{s}_{0}=\phi(0)$ is the restriction of any spinc-structure $\mathfrak{t}_{0}$ on $X$ such that $\langle c_{1}(\mathfrak{t}_{0}),[\Sigma]\rangle\equiv p\bmod 2p$.
+Since the restriction map $H^{2}(X)\rightarrow H^{2}(M)$ is reduction mod $p$, this implies that $c_{1}(\mathfrak{s}_{0})=0$, as required.
+∎
+Remark 2.4.
+As mentioned in the introduction, the line of reasoning in the proof of Proposition 2.3 can easily be extended to show that if $K$ is an almost $L$-space knot and $M$ is obtained by $p$-surgery on $K$ for some $p\geq 2g(K)-1$, then any co-orientable taut foliation on $M$ has Euler class zero.
 3. The Euler class of the normal bundle of a Seifert fibration
@@
 Denote this trivial $S^{1}$-bundle by $P\cong F\times S^{1}$.
-Fix orientations on $F$ and on the fibres so that they intersect positively.
+Fix orientations on $F$ and on the fibres so that they intersect in $M$.
 Let
@@
     is an invariant of the given Seifert fibring of $M$ called the Euler number.
+  3. (3)
+    The effect on $b,a_{1},\ldots,a_{n}$ and $e(M)$ of reversing the orientation of $M$ is to replace them by their negatives.
 Remark 3.2.
@@
 $0\to H^{1}(B)\xrightarrow{}H^{1}(M)\to\mathbb{Z}\xrightarrow{\;\delta\;}H^{2}(B)\xrightarrow{\;f^{*}\;}H^{2}(M)$
-It follows from [8, Theorem 4] that the image of 1-1 under the transgression homomorphism $\delta$ is the Euler class $bu$ of the circle bundle.
+It follows from [13, Theorem 4] that the image of 1-1 under the transgression homomorphism $\delta$ is the Euler class $bu$ of the circle bundle.
 Hence $\mbox{kernel}(f^{*})$ is generated by $bu$.
-Then by (2.1.2), $e(\nu_{M})=0$ if and only $e(\tau_{B})=\pm\chi(B)u$ is a multiple of $b$, which is condition (2.1.1).
+Then by (3.1.2), $e(\nu_{M})=0$ if and only $e(\tau_{B})=\pm\chi(B)u$ is a multiple of $bu$, which is condition (3.1.1).
 We consider the following two special cases:
@@
 By construction, the image of $c_{\sigma}]$ under $H^{2}(M,\sqcup T_{i})\to H^{2}(M)$ is the Euler class $e(\nu_{M})$.
-(See [9, Section 2] for instance.)
+(See [14, Section 2] for instance.)
     To write $c_{\sigma}]$ as a linear combination of $\alpha_{r}^{*}\otimes S^{*}$, $\beta_{r}^{*}\otimes S^{*}$, $c_{j}^{*}\otimes S^{*}$, $F^{*}$, and $D_{i}^{*}$, it suffices to compute its evaluation on the dual homology classes.
@@
 4.1. The second cohomology group of $\Gamma$
-Jankins and Neuman determined a canonical isomorphism between $H^{2}(\Gamma)$ and the abelian group having generators $A_{0},A_{1},\ldots,A_{n}$ and relations $A_{0}=b_{i}A_{i}$ for $\leq i\leq n [11$].
+Jankins and Neuman determined a canonical isomorphism between $H^{2}(\Gamma)$ and the abelian group having generators $A_{0},A_{1},\ldots,A_{n}$ and relations $A_{0}=b_{i}A_{i}$ for $1\leq i\leq n$ [16].
 (The case that $g=0$ is discussed in [16] and the general case in [15].) Each element of $H^{2}(\Gamma)$ can be uniquely expressed in the form
@@
 (4.1.2) $0\to H^{1}(\Gamma)\xrightarrow{}H^{1}(M)\to\mathbb{Z}\xrightarrow{\;\delta\;}H^{2}(\Gamma)\xrightarrow{\;\alpha^{*}\;}H^{2}(M)$
-See [8, Theorem 4]. Thus,
+See [13, Theorem 4]. Thus,
 (4.1.3) $\mbox{kernel}(\alpha^{*})=\langle e(E_{M})\rangle$
@@
 The fundamental group of $M$ acts on $\mathbb{H}^{2}$ via the composition $\varphi=\bar{\varphi}\circ\alpha$, which gives rise to a diagonal action $\psi$ of $\pi_{1}(M)$ on $\widetilde{M}\times\mathbb{H}^{2}$:
-$\psi(\gamma)(x,y)=(\gamma\cdot x,\varphi(\gamma)(y))$
+$\psi(\gamma)(x,y)=(\gamma\cdot x,\varphi(\gamma)(M))$
 This action is equivariant with respect to the projection $\widetilde{M}\times\mathbb{H}^{2}\to\widetilde{M}$, so taking quotients determines a locally-trivial $\mathbb{H}^{2}$-bundle
@@
 $E(\tilde{\nu}_{M})\xrightarrow{\;\mbox{{\footnotesize exp}}\;}\widetilde{M}=\mathbb{H}^{2}\times\mathbb{R}\xrightarrow{\;\mbox{{\footnotesize projection}}\;}\mathbb{H}^{2}$
-Now proceed as in the proof of [2, Lemma A.3], starting from its third paragraph, to show that the $\mathbb{H}^{2}$-bundle $M\times_{\psi}\mathbb{H}^{2}\to M$ is topologically equivalent to the normal bundle $\nu_{M}$ and conclude, as in [2, Proposition A4], that $e(\nu_{M})=e(\varphi)=e(\bar{\varphi}\circ\alpha)$.
+Now proceed as in the proof of [3, Lemma A.3], starting from its third paragraph, to show that the $\mathbb{H}^{2}$-bundle $M\times_{\psi}\mathbb{H}^{2}\to M$ is topologically equivalent to the normal bundle $\nu_{M}$ and conclude, as in [3, Proposition A4], that $e(\nu_{M})=e(\varphi)=e(\bar{\varphi}\circ\alpha)$.
 Thus
 (4.2.1) $e(\nu_{M})=\alpha^{*}(e(\bar{\varphi}))$
-It is well-known that $e(\bar{\varphi})$, respectively $e(\varphi)$, is the obstruction to lifting the representation $\bar{\varphi}:\Gamma\to PSL_{2}(\mathbb{R})$, respectively $\varphi:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$, to $\widetilde{PSL}_{2}(\mathbb{R})$ [14, Lemma 2].
+It is well-known that $e(\bar{\varphi})$, respectively $e(\varphi)$, is the obstruction to lifting the representation $\bar{\varphi}:\Gamma\to PSL_{2}(\mathbb{R})$, respectively $\varphi:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$, to $\widetilde{PSL}_{2}(\mathbb{R})$ [19, Lemma 2].
 An expression for $e(\bar{\varphi})$ in terms of the Jankins-Neumann generators $A_{0},A_{1},\ldots,A_{n}$ of $H^{2}(\Gamma)$ can be determined following the method of §3 of [16].
@@
 We saw in (4.2.1) that that $e(\nu_{M})=\alpha^{*}(e(\bar{\varphi}))=e(\bar{\varphi}\circ\alpha)=e(\varphi)$, where $\varphi=\bar{\varphi}\circ\alpha:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$.
-Hence, if $e(\nu_{M})=0$, then $\varphi$ lifts to a homomorphism $\pi_{1}(M)\to\widetilde{PSL}_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{R})$ [14, Lemma 2].
-It follows that $\pi_{1}(M)$ is left-orderable [6, Theorem 1.1] and therefore, $M$ admits a horizontal foliation [6, Theorem 1.3].
+Hence, if $e(\nu_{M})=0$, then $\varphi$ lifts to a homomorphism $\pi_{1}(M)\to\widetilde{PSL}_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{R})$ [19, Lemma 2].
+It follows that $\pi_{1}(M)$ is left-orderable [8, Theorem 1.1] and therefore, $M$ admits a horizontal foliation [8, Theorem 1.3].
 ∎
-4. Examples and applications
+5. Seifert fibred examples
 Recall that if $M$ is a Seifert fibred rational homology $3$-sphere, then the underlying surface of the base orbifold is either $S^{2}$ or $P^{2}$.
@@
 Proof.
-That any co-oriented taut foliation on $M$ must be horizontal is [5, Corollary 7] in the case $n=3$ and [1, Proposition 6.1] in general.
+That any co-oriented taut foliation on $M$ must be horizontal is [7, Corollary 7] in the case $n=3$ and [1, Proposition 6.1] in general.
 Thus (1) holds.
@@
 Theorem 1.8 follows immediately from this lemma and Theorem 1.11.
-Example 4.2 (Dehn surgeries on the trefoil knot).
-Let $K=T(2,3)$ be the right-handed trefoil knot, and let $\mu$ and $\lambda$ denote the standard meridian and longitude of $K$.
+Proof of Theorem 1.7.
+Let $K$ be the right-handed trefoil knot, and let $\mu$ and $\lambda$ denote the standard meridian and longitude of $K$.
 There is a natural Seifert fibre structure on the exterior $X=S^{3}\setminus\nu(K)$, whose regular fibre has slope $6\mu+\lambda$ on $\partial X$.
 Hence, $p/q$-surgery on $K$ yields a small Seifert fibred space whenever $p/q\neq 6$.
-It is known that $X(p/q)$ admits co-oriented taut foliations when $p/q<1$ [16, §3].
-On the other hand, [9, Theorem 1.7] shows that the Euler class of the co-oriented taut foliation constructed by Roberts in [16] vanishes if and only if
+It is known that $K(p/q)$ admits co-orientable taut foliations when $p/q<1$ [25, §3] and hence has a left-orderable fundamental group [4].
+On the other hand, assuming that $p>0$ and $p,q$ are coprime, [14, Theorem 1.7] shows that the Euler class of the co-orientable taut foliation constructed by Roberts in [25] vanishes if and only if
 $|q|\equiv 1\pmod{p}$.
-Combining Example 4.2 with Lemma 4.1 yields Theorem 1.6.
+This completes the proof. ∎
 The following lemma is an immediate consequence of Theorem 1.11.
 Lemma 5.2.
-If $e(\nu_{M})=0$, then $d=\gcd(b_{i},b_{j})\implies a_{i}\equiv a_{j}\mbox{ {\rm(mod $d$)}}$.
+If $e(\nu_{M})=0$, then $d=\gcd(b_{i},b_{j})$ implies $a_{i}\equiv a_{j}\mbox{ {\rm(mod $d$)}}$.
 Two sets of Seifert invariants $(b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$, $(b^{\prime},\frac{a_{1}^{\prime}}{b_{1}},\ldots,\frac{a_{n}^{\prime}}{b_{n}})$ are said to be equivalent, written
@@
 if they differ by a finite sequence of moves each of which either multiplies the coordinates by 1-1 or is of the form described in Remark 3.1(1).
-Equivalent sets of Seifert invariants yield homeomorphic manifolds.
+Equivalent sets of Seifert invariants classify Seifert structures on a closed $3$-manifold up to homeomorphism.
+See Remark 3.1.
 5.1. $\mathcal{B}$ is spherical
@@
     $e(\nu_{M})=0$ for at most one $b$ with $2-n\leq b\leq\frac{2-n}{2}$.
-4.4. Small values of $|H_{1}(M)|$
-It is known that if $|H_{1}(M)|\leq 2$, then the Euler class $e(\nu_{M})$ vanishes (see [9, Proposition 2.1] for instance).
-In this section, we show that this fact can also be obtained from the computations in this article.
-We have already discussed the case $M=M(0;b)$ in Remark 2.2, so we assume that $n>0$.
-We recall the presentation of $H^{2}(\pi_{1}(\mathcal{B}))$ in §3.1, where $H^{2}(\pi_{1}(\mathcal{B}))$ is generated by $A_{0},A_{1},\ldots,A_{n}$ with relations $A_{0}=b_{i}A_{i}$ for $i=1,\ldots,n$.
-In §3, we showed that $e(\nu_{M})=\alpha^{*}(e(\overline{\varphi}))=0$, where $\overline{\varphi}\colon\pi_{1}(\mathcal{B})\to PSL(2,\mathbb{R})$ is discrete and faithful, if and only if $\exists\,m\in\mathbb{Z}$ such that $e(\overline{\varphi})=m\,e(E_{M})$.
-See (3.1.2) and (3.1.3).
-When $M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, by (3.1.1) and (3.2.2),
-$e(E_{M})=-\bigl(bA_{0}+a_{1}A_{1}+\cdots+a_{n}A_{n}\bigr),$
-and
-$e(\overline{\varphi})=-\left((2-n)A_{0}+A_{1}+A_{2}+\cdots+A_{n}\right)$
-Suppose that $m$ is an integer such that $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i})$ for $i=1,\cdots,n$. Set
-$S_{0}:=m(bA_{0}+a_{1}A_{1}+\cdots+a_{n}A_{n})=(mb+\sum_{i=1}^{n}\left\lfloor\frac{ma_{i}}{b_{i}}\right\rfloor)A_{0}+A_{1}+\cdots+A_{n}$
-Any other integer $m^{\prime}$ for which the congruences $m^{\prime}a_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n$ hold can be written $m^{\prime}=m+rl$, where $l$ is the least common multiple of $b_{1},b_{2},\ldots,b_{n}$ and $r\in\mathbb{Z}$.
-Set $m_{r}=m+rl$.
-The reader will verify that
-(4.4.1) $S_{r}:=m_{r}(bA_{0}+a_{1}A_{1}+\cdots+a_{n}A_{n})=S_{0}+rle(M)A_{0}$
-We noted after the statement of Corollary 1.3 that $e(M)=\varepsilon|H_{1}(M)|/b_{1}b_{2}\cdots b_{n}$ for some $\varepsilon\in\{\pm 1\}$. Thus from (4.4.1) we have
-(4.4.2) $S_{r}=S_{0}+\varepsilon r\frac{|H_{1}(M)|}{(b_{1}b_{2}\cdots b_{n}/l)}A_{0}$
-When $M$ is an integer homology $3$-sphere, $H^{2}(M)=0$, so $\nu_{M}$ is trivial.
-This is reflected in the fact that $l=b_{1}b_{2}\cdots b_{n}$ in this case, so we can choose $r$ so that the coefficient of $A_{0}$ in $S_{r}$ is 2n2-n.
-Similarly, if $H_{1}(M)\cong\mathbb{Z}/2$, the fact that the tangent bundle to the Seifert fibres of $M$ is a trivial line bundle combines with the method of proof of [9, Proposition 2.1] to show that $e(\nu_{M})=0$.
-We leave it to the reader to verify this using (4.4.2) and Theorem 1.1.

### v3

@@
 2010 Mathematics Subject Classification. Primary 57M25, 57M50, 57M99
-Key words: taut foliation, Knot Floer complex, self-conjugate spinc structure, Seifert fibred manifold, normal bundle, Euler class, horizontal foliation.
+Key words: taut foliation, Euler class, knot Floer complex, self-conjugate spinc structure, Seifert fibred manifold, normal bundle, horizontal foliation.
 1. Introduction
-1.1. Rational homology spheres admitting no taut foliations with vanishing Euler class
-It is predicted by the $L$-space conjecture [4, 18] that if a closed, orientable, prime $3$-manifold $M$ admits a co-orientable taut foliation, then its fundamental group is left-orderable.
-One of the main tools for proving this is to use Thurston’s universal circle action, which to each co-oriented taut foliation $\mathcal{F}$ on an oriented rational homology $3$-sphere $M$ associates a non-trivial representation
+A long-standing problem in $3$-manifold topology is to determine which classes in the second cohomology group of a closed, oriented $3$-manifold $M$ are the Euler classes of tangent bundles of co-oriented taut foliations on $M$.
+Such classes are constrained by the fact that every orientable $3$-manifold is parallelizable, which implies that the Euler class of a co-oriented plane field over $M$ is an even class (see [18, Proposition 2.3] for instance).
+If we drop the tautness condition, every even class is realised, since each co-oriented plane field on $M$ is homotopic to an integrable one [43].
+This no longer holds when one requires that the foliation be taut.
+Indeed, only finitely many homotopy classes of co-oriented plane fields can be realised as the tangent plane field of a taut foliation [25, Corollary 1.5], [12, Corollary 6.18].
+Determining which cohomology classes can occur as Euler classes of co-orientable taut foliations remains open.
+Much work on the problem has been done when $b_{1}(M)>0$.
+In this case, Gabai has shown that taut foliations always exist [11].
+Further, results of Roussarie and Thurston show that the dual Thurston norm of the Euler class of any co-oriented taut foliation is at most one [41, 44].
+Thurston also conjectured that any class in $H^{2}(M)$ of dual Thurston norm one can be realised as the Euler class of a co-oriented taut foliation.
+However, this conjecture is now known to be false, following counterexamples constructed by Yazdi and Liu [47, 28].
+At the other extreme, one may ask whether the zero class can be realised as the Euler class of a co-oriented taut foliation.
+See [47, Question 9.1 and Question 9.2] and the related discussions therein.
+Also see [10].
+In this article, we consider the case $b_{1}(M)=0$; that is, $M$ is a rational homology sphere or equivalently, $H^{2}(M)$ is finite.
+This setting differs from the case $b_{1}(M)>0$ in two fundamental ways.
+First, the existence of co-orientable taut foliations is no longer guaranteed.
+Second, the dual Thurston norm vanishes, since $H^{2}(M)$ is finite, so the usual techniques used to analyze the Euler class when $b_{1}(M)>0$ do not apply.
+Nevertheless, the zero class remains a distinguished element of $H^{2}(M)$.
+This leads naturally to the following question (cf.
+[47, Question 9.2]):
+Question 1.1.
+If a closed, orientable rational homology sphere admits a co-orientable taut foliation, does it admit one whose Euler class is zero?
+[^1] Given a co-orientable taut foliation, the Euler class of its tangent bundle is well-defined up to sign.
+So it makes sense to discuss if a co-orientable taut foliation has vanishing Euler class.
+Question 1.1 takes on further significance in the context of the $L$-space conjecture [3, 24], which predicts that if a closed, orientable, prime $3$-manifold $M$ admits a co-orientable taut foliation, then its fundamental group is left-orderable.
+The conjecture is known to hold when $b_{1}(M)>0$ [11, 7].
+On the other hand, when $M$ is a rational homology sphere, one of the main tools for proving this is to use Thurston’s universal circle action, which to each co-oriented taut foliation $\mathcal{F}$ on an oriented rational homology $3$-sphere $M$ associates a non-trivial representation
 $\rho_{\mathcal{F}}:\pi_{1}(M)\to\mbox{Homeo}_{+}(S^{1}).$
 See [9] for the details.
-If one can show that this representation lifts to the universal covering group $\mbox{Homeo}_{\mathbb{Z}}(\mathbb{R})$ of $\mbox{Homeo}_{+}(S^{1})$, then it follows from [8, Theorem 1.1] that the fundamental group of $M$ is left-orderable.
+If one can show that this representation lifts to the universal covering group $\mbox{Homeo}_{\mathbb{Z}}(\mathbb{R})$ of $\mbox{Homeo}_{+}(S^{1})$, then it follows from [7, Theorem 1.1] that the fundamental group of $M$ is left-orderable.
 However, such a lift exists if and only if the Euler class of $\rho_{\mathcal{F}}$ vanishes.
-Since this Euler class coincides with the Euler class of the tangent bundle of $\mathcal{F}$ [5], which we denote by $e(\mathcal{F})$, we are naturally led to the following question.
-Question 1.1.
-If a closed, orientable, prime $3$-manifold admits a co-orientable taut foliation, does it admit one whose Euler class is zero?
-[^1] Given a co-orientable taut foliation, the Euler class of its tangent bundle is well-defined up to sign.
-So it makes sense to discuss if a co-orientable taut foliation has vanishing Euler class.
-Although it has been expected that the answer to the question is no, we do not know of any examples verifying this. In this article, we show that such examples occur quite frequently.
-Convention: When referring to $p/q$-surgery on a knot, we assume that $p$ and $q$ are coprime integers with $p\geq 0$.
+Since this Euler class coincides with the Euler class of the tangent bundle of $\mathcal{F}$ [5], which we denote by $e(\mathcal{F})$, we are naturally led to Question 1.1
+For the reasons discussed above, Question 1.1 has been raised often over the years.
+Its answer is yes when $H^{2}(M)$ is a (possibly trivial) group consisting of 2-torsion, since the only even class is zero in this case.
+And though it would be surprising if the answer were yes in general, no examples of manifolds for which the answer is no have appeared in the literature.
+In this article, we show that such examples are abundant, and ubiquitous amongst manifolds obtained by surgery on knots in the $3$-sphere (Theorem 1.7) and amongst Seifert fibred rational homology $3$-spheres (Theorem 1.13).
+Further, many of our examples have left-orderable fundamental groups, even though no left-order on the group arises from a lift of a universal circle representation.
+Thus we ask the following question.
+Question 1.2.
+Is there a general procedure for constructing, from a closed, orientable, prime 3-manifold which admits a co-orientable taut foliation, a non-trivial action of its fundamental group on the reals by orientation-preserving homeomorphisms?
+1.1. Surgery on knots in the $3$-sphere
+The examples derived from surgery are obtained by applying recent work of Lin, who showed that if a $3$-manifold $M$ admits a co-oriented taut foliation $\mathcal{F}$, then there is a $\mathrm{spin}^{c}$-structure $\mathfrak{s}$ on $M$ whose associated plane field is $T\mathcal{F}$ and a direct $\mathbb{F}$ summand of the $\mathbb{F}[U]$-module $HF_{\mathrm{red}}(M,\mathfrak{s})$ [27].
+Convention: When referring to $p/q$-surgery on a knot, we assume that $p$ and $q$ are coprime integers with $p>0$. We denote the manifold obtained by $p/q$-surgery on a knot $K$ by $K(p/q)$.
 Theorem 1.3.
-Let $M$ be obtained by $p/q$-surgery on a knot $K$ in the $3$-sphere. If $q$ is even and |p/q|>2g(K)1|p/q|>2g(K)-1, then $M$ does not admit a co-orientable taut foliation with Euler class zero.
-Thus the existence of a knot for which some large surgery with even denominator admits a co-orientable taut foliation shows that the answer to Question 1.1 is no.
-For instance, Roberts has shown that if $K$ is a non-trivial fibred knot in the $3$-sphere, then $K(p/q)$, the manifold obtained by $p/q$-surgery on $K$, admits a co-orientable taut foliation for all $p/q$, $p/q<1$, or p/q>1p/q>-1 depending on whether the fractional Dehn twist coefficient of $K$ is zero, positive, or negative respectively [26].
-The following corollary states the case when the fractional Dehn twist coefficient is zero.
-Examples of such knots include all hyperbolic fibred alternating knots [6, 2].
+Let $K$ be a knot in the $3$-sphere. If $q$ is even and |p/q|>2g(K)1|p/q|>2g(K)-1, then $K(p/q)$ does not admit a co-orientable taut foliation with Euler class zero.
+In [18, Theorem 1.12(1)], it was shown that if $|p/q|>2g(K)$ and $p/q$ is not an integer, then any co-orientable taut foliation on $K(p/q)$ that is transverse to the core of the surgery solid torus cannot have vanishing Euler class.
+The above theorem removes the transversality condition under the assumption that $q$ is even.
+Thus by Theorem 1.3, the existence of a knot for which some large surgery with even denominator admits a co-orientable taut foliation shows that the answer to Question 1.1 is no.
+We give some examples in the corollary below.
 Corollary 1.4.
-Suppose that $K$ is a non-trivial fibred knot in the $3$-sphere whose fractional Dehn twist coefficient is zero.
+Suppose that $K$ is a non-trivial knot in the $3$-sphere which is either
+  1. (1)
+    fibred with zero fractional Dehn twist coefficient, or
+  2. (2)
+    alternating but not a $(2,q)$-torus knot, or
+  3. (3)
+    a Montesinos knot but not a $(2,q)$-torus knot or a (2,3,2n+1)(-2,3,2n+1)-pretzel knot.
 Then for all slopes $p/q$ with $q$ even and |p/q|>2g(K)1|p/q|>2g(K)-1, $K(p/q)$ admits a co-orientable taut foliation, but none with zero Euler class.
-A similar conclusion holds for many non-fibred knots.
-Indeed, the $L$-space conjecture predicts that every irreducible rational surgery on a non-fibred knot admits a co-orientable taut foliation and this is known to be true in many instances.
-For instance, Roberts and Delman proved this for non-torus alternating knots and for Montesinos knots which are not $L$-space knots.
-Examples of manifolds having left-orderable fundamental groups and admitting co-orientable taut foliations, but none with vanishing Euler class, can also be constructed, thus showing that no left-orders on the group arise from lifts of universal circle representations.
-For instance, the figure eight knot is a fibred hyperbolic knot of genus one with zero fractional Dehn twist coefficient and each rational surgery on it has a left-orderable fundamental group [27].
-Corollary 1.4.
-Suppose that $K$ is the figure eight knot.
-Then for all slopes $p/q$ with $q$ even and $|p/q|>1$, $K(p/q)$ has a left-orderable fundamental group and admits a co-orientable taut foliation, but no co-orientable taut foliation with zero Euler class.
-By way of contrast, we note that Theorem 1.2 cannot be extended to all large surgeries on a knot without some hypotheses on the slope $p/q$.
-For instance, it is known that $p$-surgery on the figure eight knot admits co-orientable taut foliations of zero Euler class for each integer $p$ [11, 14].
-More generally, we have
-Proposition 2.3.
-Let $M$ be a rational homology sphere obtained by integer surgery on a genus one knot in the $3$-sphere. Then every co-orientable taut foliation on $M$ has Euler class zero.
-Remark 1.5.
-The line of reasoning used in the proof of Proposition 2.3 can easily be extended to show that if $K$ is an almost $L$-space knot, meaning a knot with positive almost $L$-space surgeries [9], and $M$ is obtained by $p$-surgery on $K$ for some $p\geq 2g(K)-1$, then any co-orientable taut foliation on $M$ has Euler class zero.
-This shows that for knots of arbitrarily large genus, constraints must be placed on the denominators of surgery slopes for Theorem 1.2 to hold.
-Corollary 1.6.
-Suppose that $K$ is a knot of genus one in the $3$-sphere.
-If $p$ is an integer such that $K(p)$ admits a co-orientable taut foliation, then $\pi_{1}(K(p))$ is left-orderable.
-Similarly, if $K$ is an almost $L$-space knot and $p\geq 2g(K)-1$ is an integer such that $K(p)$ admits a co-orientable taut foliation, then $\pi_{1}(K(p))$ is left-orderable.
-All the examples in Corollary 1.4 are hyperbolic.
-The simplest examples of Seifert fibred manifolds in which every co-orientable taut foliation has non-zero Euler class arise from Dehn surgery on the trefoil knot.
-Theorem 1.7.
+Part (1) of Corollary 1.4 follows from Theorem 1.3 together with Roberts’ result [40], while parts (2) and (3) follow from Theorem 1.3 and work of Delman and Roberts.
+As we mentioned above, examples of manifolds having left-orderable fundamental groups and admitting co-orientable taut foliations, but none with vanishing Euler class, can also be constructed, thus showing that no left-orders on the group arise from lifts of universal circle representations.
+Here are the simplest examples.
+More examples are provided in Theorem 1.7 and Theorem 1.13 below.
+  • The figure eight knot $K$ is a fibred hyperbolic knot of genus one with zero fractional Dehn twist coefficient and each rational surgery on it has a left-orderable fundamental group [48].
+Hence, Corollary 1.4 implies that for all slopes $p/q$ with $q$ even and $|p/q|>1$, $K(p/q)$ has a left-orderable fundamental group and admits a co-orientable taut foliation, but no co-orientable taut foliation with zero Euler class.
   • Suppose that $K$ is the right-handed trefoil knot.
-Then $K(p/q)$ admits a co-orientable taut foliation and has a left-orderable fundamental group if and only if $p/q<1$.
-It admits a co-orientable taut foliation with zero Euler class if and only if $p/q<1$ and $|q|\equiv 1\pmod{p}$.
-In general, we denote a closed oriented Seifert fibred manifold $M$ with orientable base orbifold by $M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, where $b\in\mathbb{Z}$ and $b_{i}\geq 2$ for $i=1,\ldots,n$.
-See §3.1 for a detailed explanation of the notation.
-We denote the base orbifold by $\mathcal{B}$; this is a closed orientable surface $|\mathcal{B}|$ of genus $g\geq 0$ with $n\geq 0$ cone points of orders $b_{1},...,b_{n}$.
-Recall that the Euler characteristic $\chi(\mathcal{B})$ of $\mathcal{B}$ is $\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$, and the Euler number of $M$ is $e(M)=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}$.
-Theorem 1.8.
-Suppose that $M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ is a rational homology $3$-sphere which admits a co-orientable taut foliation $\mathcal{F}$.
-Then $\mathcal{F}$ has zero Euler class if and only if there exists $m\in\mathbb{Z}$ such that
-  1. (1)
-    $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n,\;\text{and}$
-  2. (2)
-    $me(M)=\chi(\mathcal{B})$.
-We can also produce toroidal examples which are not Seifert fibred.
+Then $K(p/q)$ admits a co-orientable taut foliation (equivalently, has a left-orderable fundamental group [7]) if and only if $p/q<1$ [23, 32].
+It follows from Lemma 3.5 and [18, Theorem 1.7] that $K(p/q)$ admits a co-orientable taut foliation with zero Euler class if and only if $p/q<1$ and $|q|\equiv 1\pmod{p}$.
+All the examples in the first family are hyperbolic, while those in the second are Seifert fibred. We can also use Theorem 1.3 to produce toroidal examples which are not Seifert fibred.
 Define the JSJ graph of a knot $K$ to be the rooted tree dual to the JSJ tori of the exterior $X(K)$ of $K$, where the root vertex corresponds to the piece of the JSJ decomposition containing $\partial X(K)$.
@@
 Fibred composite knots satisfy the corollary’s hypothesis. Taking $K$ to be a connected sum of torus knots yields examples which are graph manifolds.
-1.2. Euler class of the normal bundle of a Seifert fibration
-While Theorem 1.2 is proved using the knot Floer complex, the key fact in analyzing the Euler classes of co-oriented taut foliations on a Seifert fibred rational homology $3$-sphere $M$ is that, up to isotopy, all such foliations are horizontal; that is, transverse to the Seifert fibres of $M$ (Lemma 5.1).
-Hence, the tangent bundle of $\mathcal{F}$ is isomorphic to the normal bundle $\nu_{M}$ to the Seifert fibres of $M$, while the latter makes sense even without the existence of a horizontal foliation.
-We will refer to $\nu_{M}$ as the normal bundle of the Seifert fibration. In §3 and §4, we determine the following necessary and sufficient conditions for its Euler class to vanish.
+If $K$ is an $L$-space knot then for slopes $p/q\geq 2g(K)-1$ the manifold $K(p/q)$ is an $L$-space and hence does not admit a co-orientable taut foliation.
+On the other hand, since $L$-space knots are fibred, Roberts’ work [40] shows that $K(p/q)$ has a co-orientable taut foliation if $p/q<1$.
+We establish conditions under which negative surgeries on an $L$-space knot cannot admit a co-orientable taut foliation with Euler class zero.
+These conditions are phrased in terms of congruence relations involving the torsion coefficients of the Alexander polynomial (Theorem 2.8 and Corollary 2.9), as well as the bridge index (Corollary 2.11).
+By performing negative surgeries on the (2,3,2n+1)(-2,3,2n+1)-pretzel knots, Theorem 2.8 provides the following examples.
 Theorem 1.7.
-Given a closed oriented Seifert fibred manifold $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes if and only if there exists $m\in\mathbb{Z}$ such that
+For any integer $p\geq 7$ there are infinitely many hyperbolic 3-manifolds $M$ such that $H^{2}(M)\cong\mathbb{Z}/p$ and $M$ admits a co-orientable taut foliation, but none with Euler class zero.
+Further, $\pi_{1}(M)$ is left-orderable.
+By way of contrast, note that Theorem 1.3 cannot be extended to large surgeries on all knots without some hypotheses on the slope $p/q$.
+For instance, if $M$ is a rational homology sphere for which there is a unique $\mathrm{spin}^{c}$-structure $\mathfrak{s}_{0}$ such that $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})$ is non-zero, then any co-orientable taut foliation on $M$ has Euler class zero (see Proposition 2.4).
+As a consequence, we deduce,
+Proposition 2.6.
+Let $M$ be a rational homology sphere obtained by $p$-surgery on a knot $K$ is the $3$-sphere, for some $p\in\mathbb{Z}$.
+If $K$ has genus one, or $K$ is an almost $L$-space knot  [^2] KK is an almost $L$-space knot if it has a positive almost $L$-space surgery of slope $p/q\geq 2g(K)-1$ [8].
+and $p\geq 2g(K)-1$, then any co-orientable taut foliation on $M$ has Euler class zero.
+Finally, recalling the theorem of Hu from [18] mentioned above (also see Corollary 2.9 and Remark 2.10), we ask the following question.
+Question 1.8.
+In Theorem 1.3, can the assumption that $q$ be even be replaced by the assumption $|q|>1$?
+1.2. Seifert fibred manifolds
+We denote a closed oriented Seifert fibred manifold $M$ with orientable base orbifold by $M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, where $b\in\mathbb{Z}$ and $b_{i}\geq 2$ for $i=1,\ldots,n$.
+See §3.1 for a detailed explanation of the notation.
+We denote the base orbifold by $\mathcal{B}$; this is a closed orientable surface $|\mathcal{B}|$ of genus $g\geq 0$ with $n\geq 0$ cone points of orders $b_{1},...,b_{n}$.
+Recall that the Euler characteristic $\chi(\mathcal{B})$ of $\mathcal{B}$ is $\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$, and the Euler number of $M$ is $e(M)=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}$.
+A Seifert fibred rational homology $3$-sphere $M$ which admits a co-orientable taut foliation $\mathcal{F}$ is necessarily of the form $M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ [1, Proposition 6.1].
+Further, $\mathcal{F}$ is necessarily horizontal, that is, transverse to the Seifert fibres of $M$ (see Lemma 3.5).
+As such, the tangent bundle of $\mathcal{F}$ is isomorphic to the normal bundle $\nu_{M}$ to the Seifert fibres of $M$.
+In §3 we determine necessary and sufficient conditions for the Euler class of $\nu_{M}$ to vanish (Theorem 3.3), which leads to the following theorem.
+Theorem 1.9.
+Suppose that $M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ is a rational homology $3$-sphere which admits a co-orientable taut foliation $\mathcal{F}$.
+Then $\mathcal{F}$ has zero Euler class if and only if there exists $m\in\mathbb{Z}$ such that
   1. (1)
@@
     $me(M)=\chi(\mathcal{B})$.
-Theorem 1.8 follows from Theorem 1.11 and Lemma 5.1.
-Next, we deduce the following easily verifiable necessary condition for $\nu_{M}$ to be a trivial bundle:
 Corollary 1.10.
-Given a closed oriented Seifert fibred manifold $M$, if the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes, then either $M$ is an orientable $S^{1}$-bundle over the torus or one of the following holds:
+Given a closed oriented Seifert fibred manifold $M$ which admits a co-orientable taut foliation whose Euler class vanishes, then one of the following holds:
+  1. (1)
+    $M$ is an orientable $S^{1}$-bundle over the torus,
   2. (2)
@@
     $\frac{\chi(\mathcal{B})}{e(M)}\in\mathbb{Z}.$
-Note that if $M$ is an orientable $S^{1}$-bundle over the torus, then its normal bundle is always trivial (see Remark 3.2).
-By abelianizing the presentation (4.0.1) of $\pi_{1}(M)$ obtained by viewing $M$ as doing a Dehn filling of a trivial $S^{1}$-bundle over an $n$-punctured orientable surface of genus $g$ as in §3.1, one can show that if $T_{1}(M)$ denotes the torsion subgroup of $H_{1}(M)$, then
+By abelianizing the presentation (4.0.1) of $\pi_{1}(M)$ obtained by viewing $M$ as doing a Dehn filling of a trivial $S^{1}$-bundle over an $n$-punctured orientable surface of genus $g$ as in §3.1, one can show that if $e(M)\neq 0$ and $T_{1}(M)$ denotes the torsion subgroup of $H_{1}(M)$, then
 $|T_{1}(M)|=b_{1}\cdots b_{n}|e(M)|,$
-where we take the convention that $b_{1}b_{2}\cdots b_{n}=1$ when $n=0$. Thus, it follows from Corollary 1.12 that:
+where we take the convention that $b_{1}b_{2}\cdots b_{n}=1$ when $n=0$. Thus, it follows from Theorem 1.9 that:
 Corollary 1.11.
-Consider the Seifert fibred manifold $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ and suppose that the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes.
-Then the integer $b_{1}\ldots b_{n}\chi(\mathcal{B})$ is a multiple of $|T_{1}(M)|$.
-Hence, if $M$ is a rational homology $3$-sphere, then $b_{1}\ldots b_{n}\chi(\mathcal{B})$ is a multiple of $|H_{1}(M)|$.
-We will present more examples of rational homology Seifert fibred manifolds that admit co-orientable taut foliations, but none with vanishing Euler class, in §5.3 (Examples 5.3).
+Suppose that $M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ is a rational homology $3$-sphere which admits a co-orientable taut foliation $\mathcal{F}$ with zero Euler class.
+If $e(M)\neq 0$ then the integer $b_{1}\ldots b_{n}\chi(\mathcal{B})$ is an integer multiple of $|H_{1}(M)|$.
+When the base orbifold $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ is hyperbolic, we provide a second proof of Theorem 1.9 from the viewpoint of discrete faithful representations of Fuchsian groups.
+Here is an interesting by-product of the analysis.
+Theorem 1.12.
+Let $M$ be a closed oriented Seifert fibred manifold with orientable base orbifold $\mathcal{B}$.
+If $\mathcal{B}$ is hyperbolic or Euclidean, and $e(\nu_{M})=0$, then $M$ admits a horizontal foliation.
+As an application of Theorem 1.9 we show that such examples are actually abundant in the presence of higher order torsion.
+Theorem 1.13.
+Let $G$ be a finite abelian group containing an element of order at least three.
+Then there exist infinitely many Seifert fibred manifolds $M$ such that $H_{1}(M)\cong G$ and $M$ admits a co-orientable taut foliation, but none with Euler class zero.
+Further, $\pi_{1}(M)$ is left orderable.
+Remark 1.14.
+It is known that Seifert fibred manifolds satisfy the $L$-space conjecture [7, 30, 3], so all such manifolds which admit co-orientable taut foliations have left-orderable fundamental groups even though it is often the case that this cannot be shown using universal circle representations.
+One uses, rather, the fact that co-orientable taut foliations on Seifert fibred rational homology spheres are $\mathbb{R}$-covered.
+We will present more examples of Seifert fibred manifolds that admit co-orientable taut foliations, but none with vanishing Euler class, in §5.3.
 In fact, our computations in §5 suggest that most closed Seifert fibred manifolds have nontrivial normal bundle, including the case when the base orbifold is Euclidean (i.e.
@@
 This may seem counterintuitive, since we have pointed out that the normal bundle is always trivial when the base $\mathcal{B}$ is the torus.
-We close with an interesting by-product of our analysis.
-Theorem 1.14.
-Let $M$ be a closed oriented Seifert fibred manifold with orientable base orbifold $\mathcal{B}$.
-If $\mathcal{B}$ is hyperbolic or Euclidean, and $e(\nu_{M})=0$, then $M$ admits a horizontal foliation.
-In contrast, closed Seifert fibred manifolds with spherical base orbifolds have finite fundamental groups and hence cannot admit taut foliations.
-Yet, in §5.1, we will show examples of such manifolds for which $e(\nu_{M})$ vanishes.
+In contrast, closed Seifert fibred rational homology spheres with spherical base orbifolds have finite fundamental groups and hence cannot admit taut foliations.
+Yet, in §5.1, we will construct examples of such manifolds for which $e(\nu_{M})$ vanishes.
 Plan of the paper
-We prove Theorem 1.2 and Proposition 2.3 in §2.
-In §3 we explain our notation for Seifert manifolds and prove Theorem 1.11, which gives necessary and sufficient conditions for the Euler class of the normal bundle of a Seifert fibration to vanish.
-We also prove Corollary 1.12 in this section.
-In §4 we present a different perspective on Theorem 1.11 and Corollary 1.12 in the case that the base orbifold is hyperbolic through the use of discrete faithful $PSL(2,\mathbb{R})$-representations of $\pi_{1}(\mathcal{B})$.
-As a consequence of this approach, we prove the hyperbolic case of Theorem 1.14.
-Finally, in §5 we prove Theorems 1.7 and 1.8, and provide examples of Seifert fibred manifolds whose normal bundles are (non-)trivial.
+In §2, we apply Heegaard Floer homology to produce hyperbolic $3$-manifolds obtained by Dehn filling on knots in the $3$-sphere that admit co-orientable taut foliations, but none with vanishing Euler class.
+All results stated in §1.1 are proved in this section.
+In §3 we explain our notation for Seifert manifolds and prove Theorem 3.3, which characterizes those Seifert manifolds for which normal bundle $\nu_{M}$ to the Seifert fibres is trivial.
+Theorem 1.9 then follows from this and Lemma 3.5.
+In §4 we study the Euler class of the normal bundle $\nu_{M}$ through the use of discrete faithful $PSL(2,\mathbb{R})$-representations of $\pi_{1}(\mathcal{B})$ in the case that the base orbifold $\mathcal{B}$ is hyperbolic.
+As a consequence of this approach, we provide a different proof of Theorem 3.3 when the base is hyperbolic and prove the hyperbolic case of Theorem 1.12.
+Finally, in §5 we provide examples of Seifert fibred manifolds whose normal bundles are (non-)trivial.
 We classify which occurs when $\mathcal{B}$ is spherical in §5.1 and is Euclidean in §5.2.
@@
 2. Examples from surgeries on knots in the $3$-sphere
-Theorem 1.2 is an immediate consequence of the following two lemmas. The first follows from Ozsváth and Szabó’s proof that a manifold with a co-orientable taut foliation is not an $L$-space [22].
+In this section we prove our results on surgeries on knots in the $3$-sphere.
+2.1. $\mathrm{Spin}^{c}$ structures and co-orientable taut foliations
+The set $\mathrm{Spin}^{c}(M)$ of $\mathrm{spin}^{c}$ structures $\mathfrak{s}$ on a rational homology $3$-sphere $M$ corresponds to certain classes of pairs $(\xi,\mathfrak{s}_{0})$, where $\xi$ is an oriented $2$-plane sub-bundle of the tangent bundle of $M$ and $\mathfrak{s}_{0}$ is a spin structure on $M$ [15, proof of Proposition 1].
+There is a faithful, transitive action of $H^{2}(M)$ on $\mathrm{Spin}^{c}(M)$ [15, Proposition 1].
+The Chern class of a $\mathrm{spin}^{c}$ structure $\mathfrak{s}$, denoted $c_{1}(\mathfrak{s})$, is the Euler class of its oriented plane bundle.
+The set $\mathrm{Spin}(M)$ of $\mathrm{spin}$-structures on $M$ is identified with $\{\xi_{0}\}\times\mathrm{Spin}(M)\subset\mathrm{Spin}^{c}(M)$, where $\xi_{0}$ is a trivial bundle.
+This is precisely the set of $\mathrm{spin}^{c}$ structures $\mathfrak{s}$ satisfying $c_{1}(\mathfrak{s})=0$ [15, Remark (c) page 49].
+Reversing the orientation of $\xi$ determines an involution on $\mathrm{Spin}^{c}(M)$, called conjugation, whose fixed point set is $\mathrm{Spin}(M)$ [15, Remark (c) page 49].
+Thus there are precisely $|\mathrm{Spin}(M)|=|H_{1}(M;\mathbb{Z}/2)|$ self-conjugate spinc-structures on $M$.
+We record these remarks in a lemma.
 Lemma 2.1.
-Let $M$ be a rational homology $3$-sphere admitting a co-orientable taut foliation $\mathcal{F}$.
-Then $M$ has a $\mathrm{spin}^{c}$-structure $\mathfrak{s}$ such that $HF_{\mathrm{red}}(M,\mathfrak{s})\neq 0$ and $c_{1}(\mathfrak{s})=e(\mathcal{F})$.
+Let $M$ be a rational homology $3$-sphere.
+Then there are $|H_{1}(M)|$ spinc-structures on $M$ and $|H_{1}(M;\mathbb{Z}/2)|$ self-conjugate spinc-structures.
+Further, a spinc-structure $\mathfrak{s}$ is self-conjugate if and only if $c_{1}(\mathfrak{s})=0$.
+In what follows, we take $\mathbb{F}$ to be $\mathbb{Z}/2$.
+Lemma 2.2.
+Let $M$ be a rational homology $3$-sphere admitting a co-oriented taut foliation $\mathcal{F}$.
+Then $M$ has a $\mathrm{spin}^{c}$ structure $\mathfrak{s}$ such that $HF_{\mathrm{red}}(M,\mathfrak{s})$ contains a direct $\mathbb{F}$ summand and $c_{1}(\mathfrak{s})=e(\mathcal{F})$.
 Proof.
-Let $M$ be a rational homology $3$-sphere that admits a co-orientable taut foliation $\mathcal{F}$.
+Let $M$ be a rational homology $3$-sphere that admits a co-oriented taut foliation $\mathcal{F}$.
 The tangent plane field of $\mathcal{F}$ can be perturbed to obtain a tight contact structure $\xi$ whose contact invariant $c(\xi)\neq 0\in HF_{\mathrm{red}}(M,\mathfrak{s})$, with $c_{1}(\mathfrak{s})=e(\mathcal{F})$ [36].
+Moreover, Lin deduces that the $\mathbb{F}[U]$-module $HF_{\mathrm{red}}(M,\mathfrak{s})$ contains direct $\mathbb{F}$ summand [27].
 ∎
-The second is an application of the mapping cone formula for calculating the Heegaard Floer homology of surgeries on a knot in the $3$-sphere [23, 24].
-In conjunction with Lemma 2.1, this implies Theorem 1.2 since for a 3-manifold $M$ for which $H_{1}(M)$ has odd order, the unique spinc-structure satisfying $c_{1}(\mathfrak{s})=0$ is also the unique self-conjugate spinc-structure on $M$.
+Combined, these two lemmas yield the following obstruction to the existence of taut foliations of Euler class zero and non-zero.
+Proposition 2.3.
+Let $M$ be a rational homology sphere with a co-oriented taut foliation $\mathcal{F}$.
+  1. (1)
+    If $e(\mathcal{F})=0$, then there is a self-conjugate $\mathrm{spin}^{c}$ structure $\mathfrak{s}$ such that $HF_{\mathrm{red}}(M,\mathfrak{s})$ contains a direct $\mathbb{F}$ summand.
+  2. (2)
+    If $e(\mathcal{F})\neq 0$, then there is $\mathrm{spin}^{c}$ structure $\mathfrak{s}$ such that $HF_{\mathrm{red}}(M,\mathfrak{s})$ contains a direct $\mathbb{F}$ summand and $\mathfrak{s}$ is not self-conjugate.
+In particular, Heegaard Floer homology can be used to show that certain 3-manifolds, including almost $L$-spaces, do not admit co-orientable taut foliations with non-zero Euler class.
+Proposition 2.4.
+Let $M$ be a rational homology sphere for which there is a unique $\mathrm{spin}^{c}$ structure $\mathfrak{s}_{0}$ for which $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})$ is non-zero.
+Then any co-orientable taut foliation on $M$ has Euler class zero.
+Proof.
+Since conjugation of spinc-structures induces an isomorphism on Heegaard Floer homology, we see that $\mathfrak{s}_{0}$ is necessarily self-conjugate.
+Thus Proposition 2.3 implies that $M$ cannot admit a co-orientable taut foliation with non-zero Euler class.
+∎
+2.2. Surgery on knots
+In order to apply Proposition 2.3, we need examples of manifolds for which we can understand the Heegaard Floer homology.
+For surgery on knots $K$ in the $3$-sphere, the Heegaard Floer homology can be calculated using the mapping cone formula [38, 39].
+Let $CFK^{\infty}(K)=\bigoplus_{i,j\in\mathbb{Z}}C\{(i,j)\}$ be the knot Floer complex of $K$. We will use $\bm{A}_{{\rm red},\,k}(K)$ to denotes the reduced part of
+$\bm{A}^{+}_{k}=H_{*}(C\{i\geq 0\text{ or }j\geq k\}).$
+The mapping cone formula expresses the reduced Heegaard Floer homology of surgeries on $K$ in terms of the $\bm{A}_{{\rm red},\,k}(K)$ [38, 39].
+For our purposes, the following calculation will be sufficient.
 Lemma 2.5.
-Let $M$ be obtained by $p/q$-surgery on a knot $K$ in the $3$-sphere, where $q$ is even and |p/q|>2g(K)1|p/q|>2g(K)-1.
-Then the unique self-conjugate $\mathrm{spin}^{c}$-structure $\mathfrak{s}$ on $M$ satisfies $HF_{\mathrm{red}}(M,\mathfrak{s})=0$.
+Let $M$ be obtained by $p/q$-surgery on a knot $K$ of genus $g$ in $S^{3}$ with $p/q\geq 2g-1$. Then there is an affine map
+$\phi:\mathbb{Z}/p\to\mathrm{Spin}^{c}(M),$
+such that for $i$ in the range $\frac{q-p}{2}\leq i\leq\frac{p+q-1}{2}$
+$HF_{\mathrm{red}}(M,\phi(i))\cong\begin{cases}\bm{A}_{{\rm red},\,\left\lfloor\frac{i}{q}\right\rfloor}(K),&q(1-g)\leq i\leq qg-1,\\ 0,&\text{otherwise}.\end{cases}$
 Proof.
-Without loss of generality we may assume that $p/q>0$.
-Let $CFK^{\infty}(K)=\bigoplus_{i,j\in\mathbb{Z}}C\{(i,j)\}$ be the knot Floer complex for $K$.
-Since |p/q|>2g(K)1|p/q|>2g(K)-1, the mapping cone formula for computing $HF^{+}(M)$ in terms of $CFK^{\infty}(K)$ implies that there is a labelling
+Since $p/q\geq 2g-1$, the mapping cone formula for computing $HF^{+}(M)$ in terms of $CFK^{\infty}(K)$ implies that there is a labelling
 $\phi:\mathbb{Z}/p\to\mathrm{Spin}^{c}(M),$
-such that [21, Proposition 3.5]
-$HF_{\mathrm{red}}(M,\phi(i))\cong\bigoplus_{k\in\mathbb{Z}}\bm{A}_{{\rm red},\,\lfloor\frac{pk+i}{q}\rfloor}(K),$
-where $\bm{A}_{{\rm red},\,k}(K)$ denotes the reduced part of
-$\bm{A}^{+}_{k}=H_{*}(C\{i\geq 0\text{ or }j\geq k\}).$
-However, $\bm{A}_{{\rm red},\,k}(K)=0$ for |k|>g(K)1|k|>g(K)-1 [21, Lemma 3.2]. Thus, for $0\leq i\leq p-1$, this becomes
-$HF_{\mathrm{red}}(M,\phi(i))\cong\begin{cases}\bm{A}_{{\rm red},\,\lfloor\frac{i}{q}\rfloor}(K),&0\leq i<qg(K)-q,\\ \bm{A}_{{\rm red},\,\lfloor\frac{i-p}{q}\rfloor}(K),&p-qg(K)<i<p,\\ 0,&qg(K)-q\leq i\leq p-qg(K).\end{cases}$
-In particular, for each $|j|\leq g(K)-1$, there are precisely $q$ $\mathrm{spin}^{c}$-structures for which $HF_{\mathrm{red}}(M,\mathfrak{s})$ is isomorphic to $\bm{A}_{{\rm red},\,j}(K)$, and for all other $\mathrm{spin}^{c}$-structures, $HF_{\mathrm{red}}(M,\mathfrak{s})=0$.
-Thus the set
-$S=\{\mathfrak{s}\in\mathrm{Spin}^{c}(M)\mid HF_{\mathrm{red}}(M,\mathfrak{s})\neq 0\}$
-has cardinality divisible by $q$.
-In particular, $|S|$ is even.
-Since conjugation of $\mathrm{spin}^{c}$-structures induces an isomorphism of Heegaard Floer homology, conjugation acts as an involution on $S$.
-This implies that there are an even number of self-conjugate $\mathrm{spin}^{c}$-structures in $S$.
-Since $p$ is odd, there is precisely one self-conjugate spinc-structure on $M$, so it follows that $S$ does not contain the self-conjugate spinc-structure on MM.
-[^2] An alternative line of argument is to use the fact that conjugation of spinc-structures corresponds to the involution $i\mapsto p+q-1-i$ (mod $p$) on $\mathbb{Z}/p$.
-For $q$ even, this implies that the self-conjugate spinc-structure is labelled by $s=\frac{p+q-1}{2}$.
-However since |p/q|>2g(K)1|p/q|>2g(K)-1, the above calculation shows that $HF_{\mathrm{red}}(M,\phi(s))=0$.
+such that [35, Proposition 3.5]
+$HF_{\mathrm{red}}(M,\phi(i))\cong\bigoplus_{k\in\mathbb{Z}}\bm{A}_{{\rm red},\,\lfloor\frac{pk+i}{q}\rfloor}(K).$
+Each value of $i$ (mod $p$) is represented for $i$ in the range $\frac{q-p}{2}\leq i\leq\frac{p+q-1}{2}$, so as $\bm{A}_{{\rm red},\,j}(K)=0$ for $|j|\geq g$ [35, Lemma 3.2], we have
+$HF_{\mathrm{red}}(M,\phi(i))\cong\begin{cases}\bm{A}_{{\rm red},\,\lfloor\frac{i}{q}\rfloor}(K),&q(1-g)\leq i\leq qg-1,\\ 0&\text{otherwise}.\end{cases}$
+Combined, Proposition 2.3 and Lemma 2.5 quickly yield Theorem 1.3.
+Proof of Theorem 1.3.
+Since $p$ is odd, there is an odd number of spinc-structures on $M$ of which exactly one, call it $\mathfrak{s}_{0}$, is self-conjugate (Lemma 2.1).
+After possibly reversing orientation on $M$ and replacing $K$ by its mirror image, we may assume that $p/q\geq 2g(K)-1$.
+By Lemma 2.5, we see that for each $j$ such that $|j|<g(K)$, there are precisely $q$ spinc-structures for which $HF_{\mathrm{red}}(M,\mathfrak{s})\cong\bm{A}_{{\rm red},\,j}(K)$.
+Thus the number of spinc-structures for which $HF_{\mathrm{red}}(M,\mathfrak{s})\neq 0$ is divisible by $q$ and hence even.
+Consequently, we see that there are an odd number of spinc-structures such that $HF_{\mathrm{red}}(M,\mathfrak{s})=0$.
+Since conjugation of $\mathrm{spin}^{c}$ structures induces an isomorphism of Heegaard Floer homology, this implies that $\mathfrak{s}_{0}$ on $M$ satisfies $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})=0$.
+By Proposition 2.3, this implies that no co-orientable taut foliation on $M$ can have Euler class zero.
+∎
 In the opposite direction, we note that Theorem 1.3 cannot be extended to all large surgeries on a knot $K$ without some hypotheses on the slope $p/q$.
 Proposition 2.6.
-Let $M$ be a rational homology sphere obtained by integer surgery on a genus one knot in the 3-sphere. Then every co-orientable taut foliation on $M$ has Euler class zero.
+Let $M$ be a rational homology sphere obtained by $p$-surgery on a knot $K$ is the $3$-sphere, for some $p\in\mathbb{Z}$.
+If $K$ has genus one, or $K$ is an almost $L$-space knot and $p\geq 2g(K)-1$, then any co-orientable taut foliation on $M$ has Euler class zero.
 Proof.
-Let $K$ be a genus one knot and let $M$ be obtained by $p$-surgery on $K$.
+First suppose that $K$ is an almost $L$-space knot.
+Then $K(p)$ is an almost $L$-space for each integer $p\geq 2g(K)-1$ [8] and as such, there is a unique $\mathrm{spin}^{c}$ structure $\mathfrak{s}_{0}$ for which $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\neq 0$.
+Hence, Proposition 2.4 implies that every co-orientable taut foliation on $M$ has Euler class zero.
+Next suppose that $K$ is a genus one knot and let $M$ be obtained by $p$-surgery on $K$.
 Without loss of generality we may assume that $p>0$.
-The proposition is deduced from Lemma 2.1 by showing that there is at most one spinc-structure $\mathfrak{s}_{0}$ for which $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\neq 0$ and then verifying that $c_{1}(\mathfrak{s}_{0})=0$.
-First we compute $HF_{\mathrm{red}}(M)$. For integer surgeries of slope $p\geq 2g(K)-1$, the mapping cone formula gives a labelling
-(2.0.1) $\phi:\mathbb{Z}/p\to\mathrm{Spin}^{c}(M),$
-such that [21, Proposition 3.5]
-$HF_{\mathrm{red}}(M,\phi(i))\cong\bigoplus_{k\in\mathbb{Z}}\bm{A}_{{\rm red},\,i+kp}(K).$
-However, as $g(K)=1$, we have $\bm{A}_{{\rm red},\,k}(K)=0$ for $|k|>0$ [21, Lemma 3.2]. Thus, for $0\leq i\leq p-1$, this becomes
-$HF_{\mathrm{red}}(M,\phi(i))\cong\begin{cases}\bm{A}_{{\rm red},\,0}(K),&i=0,\\ 0,&0<i\leq p-1,\end{cases}$
-giving that that there is at most one spinc-structure $\mathfrak{s}_{0}$ for which $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\neq 0$ and that this is the spinc-structure corresponding to 0 under (2.0.1).
-We verify that $c_{1}(\mathfrak{s}_{0})=0$ by unpacking the labelling in (2.0.1).
-As in [23, Section 2.4] the map $\phi$ is obtained as follows.
-Take $X$ to be the trace of $p$-surgery on $K$.
-Let $\Sigma]\in H_{2}(X)$ denote that the homology class obtained by capping off the core of the 2-handle with a Seifert surface for $K$.
-If $\mathfrak{s}\in\mathrm{Spin}^{c}(M)$ is the restriction of $\mathfrak{t}\in\mathrm{Spin}^{c}(X)$, then $\phi(i)=\mathfrak{s}$, where $i$ satisfies
-$\langle c_{1}(\mathfrak{t}),[\Sigma]\rangle+p\equiv 2i\bmod 2p.$
-Thus the spinc-structure with $\mathfrak{s}_{0}=\phi(0)$ is the restriction of any spinc-structure $\mathfrak{t}_{0}$ on $X$ such that $\langle c_{1}(\mathfrak{t}_{0}),[\Sigma]\rangle\equiv p\bmod 2p$.
-Since the restriction map $H^{2}(X)\rightarrow H^{2}(M)$ is reduction mod $p$, this implies that $c_{1}(\mathfrak{s}_{0})=0$, as required.
+In particular, we have $p\geq 2g(K)-1=1$ and we may apply Lemma 2.5 to calculate $HF_{\mathrm{red}}(M)$.
+This shows that $HF_{\mathrm{red}}(M,\phi(i))\cong 0$ for all $i\not\equiv 0\;(\bmod\;p)$.
+Thus there is at most one spinc-structure $\mathfrak{s}_{0}$ with $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\neq 0$.
+By Proposition 2.4, this implies that every co-orientable taut foliation on $M$ has Euler class zero.
+∎
+2.3. Negative surgeries on $L$-space knots
+Let $K$ be a knot in the 3-sphere. Its symmetrized Alexander polynomial can be written in the form
+$\Delta_{K}(t)=\sum_{k}a_{k}t^{k},$
+where $a_{k}=a_{-k}$ for all $k$. For $i\geq 0$ we will use $t_{i}(K)$ to denote $i$th torsion coefficient
+$t_{i}(K)=\sum_{k\geq 1}ka_{i+k}.$
+In what follows we will use $\mathcal{T}(n)$ to denote the cyclic $\mathbb{F}[U]$-module
+$\mathcal{T}(n)=\frac{\mathbb{F}[U]}{(U^{n})}.$
+Lemma 2.7.
+Let $M$ be obtained by p/q-p/q-surgery on a non-trivial $L$-space knot $K$ for $p/q>0$. If $q$ is odd, then $M$ has a self-conjugate spinc-structure $\mathfrak{s}_{0}$ such that
+$HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\cong\mathcal{T}(t_{0})\oplus\bigoplus_{k\geq 1}\mathcal{T}(t_{\lfloor\frac{2pk+q-1}{2q}\rfloor})\oplus\mathcal{T}(t_{\lfloor\frac{2pk+q-1}{2q}\rfloor}).$
+If $p+q$ is odd, then $M$ has a self-conjugate spinc-structure $\mathfrak{s}_{1}$ such that
+$HF_{\mathrm{red}}(M,\mathfrak{s}_{1})\cong\bigoplus_{k\geq 0}\mathcal{T}(t_{\lfloor\frac{2pk+p+q-1}{2q}\rfloor})\oplus\mathcal{T}(t_{\lfloor\frac{2pk+p+q-1}{2q}\rfloor})$
+Proof.
+When applied to negative surgeries on $L$-space knots [13, Lemma 18], the mapping cone formula implies that there is a labelling of spinc-structures
+$\phi:\mathbb{Z}/p\rightarrow\mathrm{Spin}^{c}(M)$
+such that
+(2.3.1) $HF_{\mathrm{red}}(M,\phi(i))\cong\bigoplus_{k\in\mathbb{Z}}\mathcal{T}(t_{|\lfloor\frac{pk+i}{q}\rfloor|}),$
+where we are using that the fact that for $L$-space knots, the quantities $H_{k}$ and $V_{k}$ are determined by the torsion coefficients of the Alexander polynomial:
+$V_{k}=H_{-k}=t_{k}(K).$
+Using the identity
+$\left\lfloor\frac{pk+i}{q}\right\rfloor=-\left\lfloor\frac{q-1-i-pk}{q}\right\rfloor$
+alongside (2.3.1) shows that $HF_{\mathrm{red}}(M,\phi(i))\cong HF_{\mathrm{red}}(M,\phi(\tau(i)))$, where $\tau$ is the involution $\tau(i)=q-1-i\;(\bmod\;p)$ on $\mathbb{Z}/p$.
+The fixed points of $\tau$ are
+$\left\{\frac{q-1}{2},\frac{p+q-1}{2}\right\}\cap\mathbb{Z}.$
+If $q$ is odd, then we take $\mathfrak{s}_{0}=\phi(\frac{q-1}{2})$.
+If $p+q$ is odd, then we take $\mathfrak{s}_{1}=\phi(\frac{p+q-1}{2})$.
+Note that $\mathfrak{s}_{0}$ and $\mathfrak{s}_{1}$ are both defined if and only if $p$ is even and that precisely one of them is defined if $p$ is odd.
+For these spinc-structures (2.3.1) reduces to
+$HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\cong\mathcal{T}(t_{0})\oplus\bigoplus_{k\geq 1}\mathcal{T}(t_{\lfloor\frac{2pk+q-1}{2q}\rfloor})\oplus\mathcal{T}(t_{\lfloor\frac{2pk+q-1}{2q}\rfloor})$
+and
+$HF_{\mathrm{red}}(M,\mathfrak{s}_{1})\cong\bigoplus_{k\geq 0}\mathcal{T}(t_{\lfloor\frac{2pk+p+q-1}{2q}\rfloor})\oplus\mathcal{T}(t_{\lfloor\frac{2pk+p+q-1}{2q}\rfloor}).$
+Since $K$ is non-trivial, $t_{0}(K)\geq 1$ which implies that $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})$ contains an odd number of cyclic summands whereas $HF_{\mathrm{red}}(M,\mathfrak{s}_{1})$ contains an even number of cyclic summands.
+So $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\not\cong HF_{\mathrm{red}}(M,\mathfrak{s}_{1})$.
+Suppose that $\mathfrak{s}_{i}$ is defined for $i=0$ or $1$. It follows from the action of $\tau$ that the set $S_{i}$ of spinc-structures
+$S_{i}=\{\mathfrak{s}\in\mathrm{Spin}^{c}(M)\mid HF_{\mathrm{red}}(M,\mathfrak{s})\cong HF_{\mathrm{red}}(M,\mathfrak{s}_{i})\}$
+has odd cardinality.
+Since Heegaard Floer is invariant under conjugation, it follows that $S_{i}$ contains a self-conjugate spinc-structure.
+Since $M$ has one self-conjugate spinc-structure if $p$ is odd and two if $p$ is even, it follows that the reduced Heegaard Floer groups in the self-conjugate spinc-structures are isomorphic to $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})$ and $HF_{\mathrm{red}}(M,\mathfrak{s}_{1})$, according to their existence.
+∎
+For $K$ a non-trivial $L$-space knot of genus $g(K)$, the non-zero coefficients of $\Delta_{K}(t)$ are equal to $\pm 1$ and alternate in sign.
+That is the symmetrized Alexander polynomial can be written in the form
+$\Delta_{K}(t)=\sum_{k=-\ell}^{\ell}(-1)^{k+\ell}t^{\beta_{k}}$
+for some sequence of integers $\beta_{l}>\dots>\beta_{-l}$ such that $\beta_{k}=-\beta_{-k}$, $\beta_{\ell}=g(K)$ and $\beta_{\ell-1}=g(K)-1$.
+From such a description, one sees that the torsion coefficients form a non-increasing sequence such that $t_{g}(K)=0$ and $t_{g-1}=1$.
+It follows that if one takes $T_{1}$ to be the quantity
+$T_{1}=|\{i\mid t_{i}(K)=1\}|,$
+then $t_{i}(K)=1$ if and only if
+$g-T_{1}\leq i\leq g-1.$
+Theorem 2.8.
+Let $M$ be obtained by p/q-p/q-surgery on an $L$-space knot $K$ for $p/q>0$.
+Then $M$ admits a co-orientable taut foliation.
+However, if $M$ admits a co-orientable taut foliation with Euler class zero, then
+q(2g(K)1)=r+mpq(2g(K)-1)=r+mp
+for some integers $m\geq 0$ and $1\leq r<2qT_{1}$, where $T_{1}=|\{i\mid t_{i}(K)=1\}|$.
+Proof.
+An $L$-space knot is fibred and its monodromy has positive fractional Dehn twist coefficient [34, 16].
+Consequently Roberts [40] shows that any negative surgery on an $L$-space knot admits a co-orientable taut foliation.
+Now suppose that $M$ admits a co-orientable taut foliation with Euler class zero.
+From Lemma 2.7, Proposition 2.3 and the discussion prior to this proposition, we see that there is a non-negative integer $m$ such that pm+q1pm+q-1 is even and
+$g-T_{1}\leq\left\lfloor\frac{pm+q-1}{2q}\right\rfloor\leq g-1.$
+Since pm+q1pm+q-1 is even, $\frac{pm+q-1}{2q}$ can be expressed as a fraction with denominator $q$ and so we have
+$g-T_{1}\leq\frac{pm+q-1}{2q}\leq g-\frac{1}{q}.$
+Rearranging this implies that
+$q(2g-1)-2qT_{1}+1\leq pm\leq q(2g-1)-1.$
+This implies that there is $1\leq r\leq 2qT_{1}-1$ such that
+q(2g1)=r+mp.q(2g-1)=r+mp.
+2.4. Applications of Theorem 2.8
+Using Theorem 2.8, we can generate many examples of manifolds admitting a co-orientable taut foliation, but none with Euler class zero.
+For instance, it allows us to drop the condition that $q$ be even from Theorem 1.3 for many $L$-space knots.
+Corollary 2.9.
+Let $M$ be a manifold obtained by p/q-p/q-surgery on an $L$-space knot $K$ with $t_{0}(K)\geq 2$, where $p/q\geq 2g(K)-1$.
+Then $M$ admits a co-orientable taut foliation but none with Euler class zero.
+Proof.
+If $M$ admits a co-orientable taut foliation with zero Euler class then Theorem 2.8 implies that q(2g(K)1)=r+mpq(2g(K)-1)=r+mp for integers $m\geq 0$ and $r$, where $1\leq r<2qT_{1}$.
+If $m>0$, then p/q=(2g(K)1)/mr/qm<2g(K)1p/q=(2g(K)-1)/m-r/qm<2g(K)-1, contrary to our hypotheses.
+Thus $m=0$ and therefore $q(2g(K)-1)=r<2qT_{1}$.
+On the other hand, the condition that $t_{0}(K)\geq 2$ implies that $T_{1}\leq g(K)-1$, so q(2g(K)1)=r<2q(g(K)1)q(2g(K)-1)=r<2q(g(K)-1), which is impossible.
+Thus $M$ admits no co-orientable taut foliation with Euler class zero.
 ∎
 Remark 2.10.
-As mentioned in the introduction, the line of reasoning in the proof of Proposition 2.3 can easily be extended to show that if $K$ is an almost $L$-space knot and $M$ is obtained by $p$-surgery on $K$ for some $p\geq 2g(K)-1$, then any co-orientable taut foliation on $M$ has Euler class zero.
+The condition $t_{0}(K)\geq 2$ appears to be generic among $L$-space knots.
+The authors are aware of only five $L$-space knots for which it does not hold: the three torus knots $T(2,3)$, $T(2,5)$ and $T(3,4)$, the $(2,3)$-cable of $T(2,3)$, and the $(2,5)$-cable of $T(2,3)$, of genus $1,2,3,3,$ and $4$, respectively.
+See [26].
+Since $t_{0}(K)=1$ if and only if $\Delta_{K}(t)=(t^{g}+t^{-g})-(t^{g-1}+t^{-(g-1)})+1$, where $g=g(K)$, it follows from [42, Corollary 4.5] that if $K$ has a lens space surgery (so, conjecturally, is a Berge knot) and $g(K)\geq 4$ then $t_{0}(K)\geq 2$.
+It will be convenient to note that one can bound the quantity $T_{1}$ appearing in Theorem 2.8 using the bridge index.
+Corollary 2.11.
+Let $K$ be an $L$-space knot of bridge index $\mathrm{br}(K)$ and let $M$ be obtained by p/q-p/q-surgery on $K$ for $p/q>0$.
+Then $M$ admits a co-orientable taut foliation.
+However, if $M$ admits a co-orientable taut foliation with Euler class zero, then
+q(2g(K)1)=r+mpq(2g(K)-1)=r+mp
+for some integers $m\geq 0$ and $1\leq r<2q\mathrm{br}(K)$.
+Proof.
+Given Theorem 2.8, it suffices to prove that $T_{1}\leq\mathrm{br}(K)$.
+This follows from the properties of the invariant $\mathrm{Ord}_{v}(K)$ defined in [20].
+The symmetrized Alexander polynomial of a non-trivial $L$-space knot can be written in the form
+$\Delta_{K}(t)=t^{g}-t^{g-1}+t^{g-T_{1}}+\text{lower order terms}$
+The calculation of $\mathrm{Ord}_{v}(K)$ for $L$-space knots [20, Lemma 5.1] shows that
+$T_{1}\leq\mathrm{Ord}_{v}(K)+1.$
+However, for any knot $K$ the bound $\mathrm{Ord}_{v}(K)+1\leq\mathrm{br}(K)$ holds [20, Corollary 1.9]. ∎
+Theorem 1.7.
+For any integer $p\geq 7$ there are infinitely many hyperbolic 3-manifolds $M$ such that $H^{2}(M)\cong\mathbb{Z}/p$ and $M$ admits a co-orientable taut foliation, but none with Euler class zero.
+Further, $\pi_{1}(M)$ is left-orderable.
+Proof.
+For $g\geq 5$, let $K_{g}$ denote the (2,3,2g3)(-2,3,2g-3)-pretzel knot.
+The knot $K_{g}$ is a hyperbolic $L$-space knot [37].
+Moreover, we have $br(K_{g})=3$ and $g(K)=g$.
+Let $M$ be obtained by p/q-p/q-surgery on $K_{g}$ for $p/q>0$.
+The surgery slope 4g14g-1 is an exceptional surgery slope for $K_{g}$ since it yields a Seifert fibred space [4, Proposition 16].
+Since $\Delta(4g-1,-p/q)=|(4g-1)q+p|\geq 20$, the known bounds on exceptional fillings imply that $M$ is a hyperbolic manifold [29].
+The work of Roberts shows that there exists a co-orientable taut foliation on $M$.
+Moreover, if $g\neq 6$, then $M$ has left-orderable fundamental group [45, 46].
+On the other hand, if $M$ admits a co-orientable taut foliation with Euler class zero, then Corollary 2.11 implies
+$q(2g(K_{g})-1)\equiv r\;(\bmod\;p)$
+where $1\leq r\leq 6q-1$. Thus we see that $M$ does not admit a co-orientable taut foliation with Euler class zero, whenever $p/q\geq 7$ and $g\equiv 0\;(\bmod\;p)$. ∎
 3. The Euler class of the normal bundle of a Seifert fibration
 Let $M$ be a closed, oriented Seifert fibred manifold whose base orbifold has an orientable underlying space.
-In this section, we compute the Euler class of the normal bundle of the Seifert fibration on $M$ and determine conditions for it to vanish (Theorem 1.11 and Corollary 1.12).
+In this section, we compute the Euler class of $\nu_{M}$, the normal bundle to the Seifert fibres of $M$, and determine conditions for it to vanish.
 3.1. Seifert notation
@@
     This description of the Seifert fibred manifold is not unique.
-More precisely, we have that $M\Bigl(g;\,\tfrac{a_{0}}{b_{0}},\ldots,\tfrac{a_{n}}{b_{n}}\Bigr)$ is homeomorphic to $M\Bigl(g;\,\tfrac{a^{\prime}_{0}}{b_{0}},\ldots,\tfrac{a^{\prime}_{n}}{b_{n}}\Bigr)$ by an orientation-preserving homeomorphism preserving fibres if and only if up to reordering,
+More precisely, $M\Bigl(g;\,\tfrac{a_{0}}{b_{0}},\ldots,\tfrac{a_{n}}{b_{n}}\Bigr)$ is homeomorphic to $M\Bigl(g;\,\tfrac{a^{\prime}_{0}}{b_{0}},\ldots,\tfrac{a^{\prime}_{n}}{b_{n}}\Bigr)$ by an orientation-preserving homeomorphism preserving fibres if and only if up to reordering,
     $\frac{a_{i}}{b_{i}}-\frac{a^{\prime}_{i}}{b_{i}}\in\mathbb{Z}\quad\text{and}\quad\sum_{i=0}^{n}\frac{a_{i}}{b_{i}}=\sum_{i=0}^{n}\frac{a^{\prime}_{i}}{b_{i}}.$
@@
 $0\to H^{1}(B)\xrightarrow{}H^{1}(M)\to\mathbb{Z}\xrightarrow{\;\delta\;}H^{2}(B)\xrightarrow{\;f^{*}\;}H^{2}(M)$
-It follows from [13, Theorem 4] that the image of 1-1 under the transgression homomorphism $\delta$ is the Euler class $bu$ of the circle bundle.
+It follows from [17, Theorem 4] that the image of 1-1 under the transgression homomorphism $\delta$ is the Euler class $bu$ of the circle bundle.
 Hence $\mbox{kernel}(f^{*})$ is generated by $bu$.
@@
     When $B$ is the torus, (3.1.2) implies that the Euler class $e(\nu_{M})$ is always trivial, since the tangent bundle $T\mathcal{B}$ is trivial,
+3.2. The triviality of the normal bundle of a Seifert fibring
+Here is our main result.
+Theorem 3.3.
+Given a closed oriented Seifert fibred manifold $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes if and only if there exists $m\in\mathbb{Z}$ such that
+  1. (1)
+    $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n,\;\text{and}$
+  2. (2)
+    $me(M)=\chi(\mathcal{B})$.
+Corollary 3.4.
+Suppose that $M$ is a closed oriented Seifert fibred manifold $M$ such that the Euler class of $\nu_{M}$ vanishes. Then one of the following holds:
+  1. (1)
+    $M$ is an orientable $S^{1}$-bundle over the torus,
+  2. (2)
+    $e(M)=\chi(\mathcal{B})=0$,
+  3. (3)
+    $e(M)\neq 0,\,\chi(\mathcal{B})\neq 0,$ and
+    $\frac{\chi(\mathcal{B})}{e(M)}\in\mathbb{Z}.$
+The following lemma is key to applying these results to the Euler classes of foliations on Seifert fibred rational homology $3$-spheres.
+Lemma 3.5.
+Suppose that $M$ is a Seifert fibred rational homology $3$-sphere. Then
+$(1)$ Any co-orientable taut foliation on $M$ is horizontal. That is, it can be isotoped to be everywhere transverse to the Seifert fibres of $M$.
+$(2)$ The tangent space to any co-orientable taut foliation on $M$ is isomorphic to $\nu_{M}$ as an unoriented bundle.
+In particular, the Euler class of any co-oriented taut foliation on $M$ agrees, up to sign, with that of $\nu_{M}$.
+Proof.
+That any co-oriented taut foliation on $M$ must be horizontal is [6, Corollary 7] in the case $n=3$ and [1, Proposition 6.1] in general.
+Thus (1) holds.
+Consequently, the tangent bundle of any such foliation is isomorphic to $\nu_{M}$ (as an unoriented bundle), which implies (2).
+∎
+Theorem 1.9 follows immediately from this lemma and Theorem 3.3.
 3.3. Proof of Corollary 3.4
@@
 By construction, the image of $c_{\sigma}]$ under $H^{2}(M,\sqcup T_{i})\to H^{2}(M)$ is the Euler class $e(\nu_{M})$.
-(See [14, Section 2] for instance.)
+(See [18, Section 2] for instance.)
     To write $c_{\sigma}]$ as a linear combination of $\alpha_{r}^{*}\otimes S^{*}$, $\beta_{r}^{*}\otimes S^{*}$, $c_{j}^{*}\otimes S^{*}$, $F^{*}$, and $D_{i}^{*}$, it suffices to compute its evaluation on the dual homology classes.
@@
 This proves the theorem. ∎
-4. Discrete representations of Fuchsian groups and the normal bundle of a Seifert fibration
+4. Representations of Fuchsian groups and the normal bundle of a Seifert fibration
 In this section we provide alternative proofs of Theorem 3.3 and Corollary 3.4 from the perspective of discrete faithful representations of Fuchsian groups.
@@
 (4.1.2) $0\to H^{1}(\Gamma)\xrightarrow{}H^{1}(M)\to\mathbb{Z}\xrightarrow{\;\delta\;}H^{2}(\Gamma)\xrightarrow{\;\alpha^{*}\;}H^{2}(M)$
-See [13, Theorem 4]. Thus,
+See [17, Theorem 4]. Thus,
 (4.1.3) $\mbox{kernel}(\alpha^{*})=\langle e(E_{M})\rangle$
@@
 $E(\tilde{\nu}_{M})\xrightarrow{\;\mbox{{\footnotesize exp}}\;}\widetilde{M}=\mathbb{H}^{2}\times\mathbb{R}\xrightarrow{\;\mbox{{\footnotesize projection}}\;}\mathbb{H}^{2}$
-Now proceed as in the proof of [3, Lemma A.3], starting from its third paragraph, to show that the $\mathbb{H}^{2}$-bundle $M\times_{\psi}\mathbb{H}^{2}\to M$ is topologically equivalent to the normal bundle $\nu_{M}$ and conclude, as in [3, Proposition A4], that $e(\nu_{M})=e(\varphi)=e(\bar{\varphi}\circ\alpha)$.
+Now proceed as in the proof of [2, Lemma A.3], starting from its third paragraph, to show that the $\mathbb{H}^{2}$-bundle $M\times_{\psi}\mathbb{H}^{2}\to M$ is topologically equivalent to the normal bundle $\nu_{M}$ and conclude, as in [2, Proposition A4], that $e(\nu_{M})=e(\varphi)=e(\bar{\varphi}\circ\alpha)$.
 Thus
 (4.2.1) $e(\nu_{M})=\alpha^{*}(e(\bar{\varphi}))$
-It is well-known that $e(\bar{\varphi})$, respectively $e(\varphi)$, is the obstruction to lifting the representation $\bar{\varphi}:\Gamma\to PSL_{2}(\mathbb{R})$, respectively $\varphi:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$, to $\widetilde{PSL}_{2}(\mathbb{R})$ [19, Lemma 2].
+It is well-known that $e(\bar{\varphi})$, respectively $e(\varphi)$, is the obstruction to lifting the representation $\bar{\varphi}:\Gamma\to PSL_{2}(\mathbb{R})$, respectively $\varphi:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$, to $\widetilde{PSL}_{2}(\mathbb{R})$ [31, Lemma 2].
 An expression for $e(\bar{\varphi})$ in terms of the Jankins-Neumann generators $A_{0},A_{1},\ldots,A_{n}$ of $H^{2}(\Gamma)$ can be determined following the method of §3 of [22].
@@
 Then by (4.1.1),
-(4.2.4) $\theta(e(E_{M}))=-l(b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}})=-le(M),$
+(4.2.4) $\theta(e(E_{M}))=-l\left(b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}\right)=-le(M),$
 where $e(M)$ is the Euler number of $M$ (see §3.1).
 Next note that (4.2.2) implies that
-(4.2.5) $\theta(e(\bar{\varphi}))=l(n-\chi(|\mathcal{B}|)-\sum_{i}\frac{1}{b_{i}})=-l\chi(\mathcal{B})$
+(4.2.5) $\theta(e(\bar{\varphi}))=l\left(n-\chi(|\mathcal{B}|)-\sum_{i}\frac{1}{b_{i}}\right)=-l\chi(\mathcal{B})$
 Suppose that $e(\nu_{M})=0$.
@@
 We saw in (4.2.1) that that $e(\nu_{M})=\alpha^{*}(e(\bar{\varphi}))=e(\bar{\varphi}\circ\alpha)=e(\varphi)$, where $\varphi=\bar{\varphi}\circ\alpha:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$.
-Hence, if $e(\nu_{M})=0$, then $\varphi$ lifts to a homomorphism $\pi_{1}(M)\to\widetilde{PSL}_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{R})$ [19, Lemma 2].
-It follows that $\pi_{1}(M)$ is left-orderable [8, Theorem 1.1] and therefore, $M$ admits a horizontal foliation [8, Theorem 1.3].
+Hence, if $e(\nu_{M})=0$, then $\varphi$ lifts to a homomorphism $\pi_{1}(M)\to\widetilde{PSL}_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{R})$ [31, Lemma 2].
+It follows that $\pi_{1}(M)$ is left-orderable [7, Theorem 1.1] and therefore, $M$ admits a horizontal foliation [7, Theorem 1.3].
 ∎
@@
 Orient $\mathcal{B}$ and endow $\nu_{M}$, the normal bundle to the Seifert fibres of $M$, with the induced orientation.
-The following lemma is key to our applications.
-Lemma 5.1.
-Suppose that $M$ is a Seifert fibred rational homology $3$-sphere. Then
-$(1)$ Any co-orientable taut foliation on $M$ is horizontal. That is, it can be isotoped to be everywhere transverse to the Seifert fibres of $M$.
-$(2)$ The tangent space to any co-orientable taut foliation on $M$ is isomorphic to $\nu_{M}$ as an unoriented bundle.
-In particular, the Euler class of any co-oriented taut foliation on $M$ agrees, up to sign, with that of $\nu_{M}$.
+Proposition 5.1.
+Let $M$ be a Seifert fibred rational homology sphere with $n$ exceptional fibres. If $n\geq|H_{1}(M)|+3$, then $M$ admits a coorientable taut foliation.
 Proof.
-That any co-oriented taut foliation on $M$ must be horizontal is [7, Corollary 7] in the case $n=3$ and [1, Proposition 6.1] in general.
-Thus (1) holds.
-Consequently, the tangent bundle of any such foliation is isomorphic to $\nu_{M}$ (as an unoriented bundle), which implies (2).
-∎
-Theorem 1.8 follows immediately from this lemma and Theorem 1.11.
-Proof of Theorem 1.7.
-Let $K$ be the right-handed trefoil knot, and let $\mu$ and $\lambda$ denote the standard meridian and longitude of $K$.
-There is a natural Seifert fibre structure on the exterior $X=S^{3}\setminus\nu(K)$, whose regular fibre has slope $6\mu+\lambda$ on $\partial X$.
-Hence, $p/q$-surgery on $K$ yields a small Seifert fibred space whenever $p/q\neq 6$.
-It is known that $K(p/q)$ admits co-orientable taut foliations when $p/q<1$ [25, §3] and hence has a left-orderable fundamental group [4].
-On the other hand, assuming that $p>0$ and $p,q$ are coprime, [14, Theorem 1.7] shows that the Euler class of the co-orientable taut foliation constructed by Roberts in [25] vanishes if and only if
-$|q|\equiv 1\pmod{p}$.
-This completes the proof. ∎
+Suppose that $M$ can be expressed in the form
+$M=M\left(0;b,\frac{a_{1}}{b_{1}},\dots,\frac{a_{n}}{b_{n}}\right),$
+where $0<\frac{a_{1}}{b_{1}}<1$ for all $i$. We give two proofs, one from the viewpoint of Heegard Floer homology, the other from the viewpoint of foliations transverse to the Seifert fibres of $M$.
+The Heegard Floer approach. Up to reversing orientation, we can assume that $e(M)<0$.
+Suppose that $M$ does not admit a coorientable taut foliation, or equivalently, that $M$ is an $L$-space (Lemma 3.5 and [30]).
+Némethi [33, § 11] provides an algorithm for calculating the Heegaard Floer homology of $M$ in terms of its Seifert invariants.
+This algorithm shows that if $M$ is an $L$-space then the function
+$\Delta(i)=1-ib-\sum_{j=1}^{n}\left\lceil\frac{ia_{j}}{b_{j}}\right\rceil$
+satisfies $\Delta(i)\geq 0$ for all $i\geq 0$. [^4] Actually this is equivalent to being an $L$-space. In particular, one calculates that
+$\Delta(1)=1-b-n\geq 0$
+and that
+$\Delta((b_{1}\cdots b_{n})-1)=1+|H_{1}(M)|+b\geq 0.$
+Summing these two implies that $n\leq|H_{1}(M)|+2$, as required.
+The transverse foliation approach.
+After reversing orientation we have $e(M)>0$, and after reindexing we can suppose that $b_{1}\geq b_{i}$ for each $i$.
+Then with $p=b_{1}b_{2}\cdots b_{n}$ and $\sigma=\sum_{i=1}^{n}a_{i}/b_{i}$, we have $p(b+\sigma)=pe(Y)=|H_{1}(M)|$.
+Therefore, as $|H_{1}(M)|\leq n-3<nb_{2}b_{3}\cdots b_{n}$,
+$\frac{|H_{1}(M)|}{p}-n<\frac{|H_{1}(M)|}{p}-\sigma=b\leq\frac{|H_{1}(M)|}{p}-\frac{n}{b_{1}}=\frac{1}{b_{1}}\left(\frac{|H_{1}(M)|-nb_{2}b_{3}\cdots b_{n}}{b_{2}b_{3}\cdots b_{n}}\right)<0.$
+It follows that $1-n\leq b\leq-1$.
+If $2-n\leq b\leq-2$, $M$ admits a co-orientable taut foliation by [23, Theorem 2].
+If b=1b=-1, set $m=p+|H_{1}(M)|-(n-2)$. Then $2\leq m<p$ and if $p_{i}=p/b_{i}\in\mathbb{Z}$ we have
+  • $a_{1}p_{1}+a_{2}p_{2}+\cdots+a_{n}p_{n}=p\sigma=p+|H_{1}(M)|=m+n-2$;
+  • $a_{i}/b_{i}=a_{i}p_{i}/p<a_{i}p_{i}/m$.
+It follows that the $n$-tuple $(a_{1}/b_{1},\ldots,a_{n}/b_{n})$ lies in $R_{n}(m)$, where
+$R_{n}(m)=\{(\gamma_{1},\gamma_{2},\ldots,\gamma_{n})\in(0,1)^{n}\;|\;\exists\;(r_{1},r_{2},\ldots,r_{n})\in\mathbb{Z}^{n}\mbox{ such that }$
+[…]
+Corollary 3.3 of [23] now implies that $M$ admits a co-orientable taut foliation.
+A similar argument can be applied when b=1nb=1-n, after observing that
+$M=M\left(0;1-n,\frac{a_{1}}{b_{1}},\dots,\frac{a_{n}}{b_{n}}\right)$ $\cong$ $M\left(0;n-1,-\frac{a_{1}}{b_{1}},\dots,-\frac{a_{n}}{b_{n}}\right)$
+$=$ $M\left(0;-1,\frac{b_{1}-a_{1}}{b_{1}},\ldots,\frac{b_{n}-a_{n}}{b_{n}}\right)$
 The following lemma is an immediate consequence of Theorem 3.3.
@@
 If $e(\nu_{M})=0$, then $d=\gcd(b_{i},b_{j})$ implies $a_{i}\equiv a_{j}\mbox{ {\rm(mod $d$)}}$.
+Proof of Theorem 1.13.
+Let $G$ be a finite abelian group of containing an element of order at least $3$.
+We may write $G\cong\frac{\mathbb{Z}}{d_{1}}\oplus\dots\oplus\frac{\mathbb{Z}}{d_{k}}$, where $d_{i+1}$ divides $d_{i}$ for each $i$ and $d_{1}\geq 3$.
+Consider the Seifert fibred space
+(5.0.1) $M=M\left(0;0,\frac{a}{d_{1}^{2}},\frac{1}{d_{1}},\dots,\frac{1}{d_{k}}\right),$
+where $a=-1-\sum_{i=1}^{k}\frac{d_{1}^{2}}{d_{i}}$.
+For such an $a$ we have $a\equiv-1\;(\bmod\;d_{1})$.
+Consequently $a$ is coprime to $d_{1}^{2}$.
+Moreover, as $d_{1}>2$, we see from Lemma 5.2 that $e(\nu_{M})\neq 0$.
+Calculating the homology using (4.0.1) shows that $H_{1}(M)\cong G$ (compare [19, § 4]).
+Next observe that we may perform surgery on the regular fibres in order to obtain a new Seifert fibered space with the same first homology group.
+More precisely, suppose that $Y$ is the Seifert fibred rational homology $3$-sphere
+$Y=M\left(0;b,\frac{a_{1}}{b_{1}},\dots,\frac{a_{n}}{b_{n}}\right)$
+oriented so that $e(Y)>0$ and let $H=|H_{1}(Y)|$ denote the order of the first homology of $Y$. Then one can verify that the manifold
+$Y^{\prime}=M\left(0;b,\frac{a_{1}}{b_{1}},\dots,\frac{a_{n}}{b_{n}},\frac{-H^{2}}{(b_{1}\cdots b_{n})H+1}\right),$
+satisfies $H_{1}(Y)\cong H_{1}(Y^{\prime})$. First, one can calculate that
+$|H_{1}(Y^{\prime})|=(b_{1}b_{2}\cdots b_{n})(H(b_{1}b_{2}\cdots b_{n})+1)e(Y^{\prime})=|H_{1}(Y)|.$
+On the other hand, consideration of the presentations of $H_{1}(Y)$ and $H_{1}(Y^{\prime})$ obtained from (4.0.1) shows that there is an epimorphism $H_{1}(Y^{\prime})\to H_{1}(Y)$ induced by sending $x_{i}$ to $x_{i}$ for $1\leq i\leq n$, $x_{n+1}$ to $1$, and $h$ to $h$.
+Thus $H_{1}(Y^{\prime})\cong H_{1}(Y)$.
+Now starting with the manifold $M$ as defined in (5.0.1), and applying the above construction, we obtain an infinite sequence of Seifert fibred spaces, none of which admit a co-orientable taut foliation with Euler class zero by condition (1) of Theorem 3.3.
+It follows from Proposition 5.1 that if
+$Y=M\left(0;b,\frac{a_{1}}{b_{1}},\dots,\frac{a_{n}}{b_{n}}\right)$
+is a Seifert fibred rational homology sphere which is an $L$-space or, equivalently, does not admit a coorientable taut foliation, then $n\leq 2+|H_{1}(Y)|$.
+Thus, the manifolds in this sequence will eventually admit co-orientable taut foliations.
+∎
 Two sets of Seifert invariants $(b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$, $(b^{\prime},\frac{a_{1}^{\prime}}{b_{1}},\ldots,\frac{a_{n}^{\prime}}{b_{n}})$ are said to be equivalent, written
-$(b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})\sim(b^{\prime},\frac{a_{1}^{\prime}}{b_{1}},\ldots,\frac{a_{n}^{\prime}}{b_{n}})$
+$\left(b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}}\right)\sim\left(b^{\prime},\frac{a_{1}^{\prime}}{b_{1}},\ldots,\frac{a_{n}^{\prime}}{b_{n}}\right)$
 if they differ by a finite sequence of moves each of which either multiplies the coordinates by 1-1 or is of the form described in Remark 3.1(1).
-Equivalent sets of Seifert invariants classify Seifert structures on a closed $3$-manifold up to homeomorphism.
+Equivalence classes of Seifert invariants classify Seifert structures on a closed $3$-manifold up to homeomorphism.
 See Remark 3.1.
@@
 By assumption, $a_{i}=1$ when $b_{i}=2$. If $b_{2}=3$ and $a_{2}=2$, then the Seifert invariants of $M$ are
-$(b,\frac{1}{2},\frac{2}{3},\frac{a_{3}}{b_{3}})\sim(b+1,\frac{1}{2},\frac{-1}{3},\frac{a_{3}}{b_{3}})\sim(-b-1,-\frac{1}{2},\frac{1}{3},-\frac{a_{3}}{b_{3}})\sim(-b-3,\frac{1}{2},\frac{1}{3},\frac{b_{3}-a_{3}}{b_{3}})$
+$\left(b,\frac{1}{2},\frac{2}{3},\frac{a_{3}}{b_{3}}\right)\sim\left(b+1,\frac{1}{2},\frac{-1}{3},\frac{a_{3}}{b_{3}}\right)\sim\left(-b-1,-\frac{1}{2},\frac{1}{3},-\frac{a_{3}}{b_{3}}\right)\sim\left(-b-3,\frac{1}{2},\frac{1}{3},\frac{b_{3}-a_{3}}{b_{3}}\right)$
 Thus we can assume that $M$ has Seifert invariants $(b,\frac{1}{2},\frac{1}{b_{2}},\frac{a}{b_{3}})$.
@@
 It is interesting to note that the values of the Seifert invariants when $e(\nu_{M})=0$ correspond to those which realise the minimal values of $|H_{1}(M)|$ in the given families.
+Remark 5.3.
+Let $T(r,s)$ be the $(r,s)$-torus knot, where $r<s$ are relatively prime integers greater than 1, and let $g$ denote its genus.
+If $p/q\neq rs$ then $M=T(r,s)(p/q)$ has a Seifert fibration induced from that on the exterior of $T(r,s)$, and we can use Theorem 3.3 to determine whether or not the Euler class $e(\nu_{M})$ is zero.
+It turns out that if $(r,s)\neq(2,3)$, then $e(\nu_{M})=0$ if and only if $(2g-1)|q|\equiv 1(\text{{\rm mod} }p)$.
+Since $g>1$, this implies that p/q<2g1p/q<2g-1, and therefore $M$ admits a co-orientable taut foliation.
+On the other hand, for the trefoil $T(2,3)$ we get that $e(\nu_{M})=0$ if and only if $|q|\equiv 1(\text{{\rm mod} }p)$ and $p/q<6$.
+Thus either $p/q<1$, in which case $M$ has a co-orientable taut foliation, or $p/q=1,2,3,4$ or $5$.
+The latter five manifolds $M$ are, respectively, those with Seifert invariants (1,1/2,1/3,1/5),(1,1/2,1/3,1/4),(1,1/2,1/3,1/3),(1,1/2,1/2,1/3)(-1,1/2,1/3,1/5),(-1,1/2,1/3,1/4),(-1,1/2,1/3,1/3),(-1,1/2,1/2,1/3), and $L(5,1)$.
+As these have finite fundamental groups, they do not admit co-orientable taut foliations.
 5.2. $\mathcal{B}$ is Euclidean
@@
 We can assume that $a_{i}=1$ in all cases.
-This is obvious when $b_{i}=2$ and using Lemma 5.2 and the arguments of §5.1 deals with the case that some $b_{i}>2$.
+This is obvious when $b_{i}=2$ and using Lemma 5.2 and the arguments of §5.1 deal with the case that some $b_{i}>2$.
 As $\mathcal{B}$ is Euclidean, its Euler characteristic is zero.
@@
 5.3. $\mathcal{B}$ is hyperbolic
-We assume in this section that the Seifert invariants of
-$M=M(0;b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$
-have been normalised so that $0<a_{i}<b_{i}$ for $1\leq i\leq n$.
+Throughout this subsection, we assume that $\mathcal{B}=S^{2}(b_{1},...,b_{n})$ is hyperbolic.
+We also assume that $M=M(0;b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$ is a Seifert manifold with base orbifold $\mathcal{B}$ whose Seifert invariants have been normalised so that $0<a_{i}<b_{i}$ for $1\leq i\leq n$.
+Remark 5.4.
 The results of [23] and [32] show that $M$ admits a horizontal foliation if and only if one of the following conditions holds:
@@
     b=1nb=1-n and there are coprime integers $0<c<d$ such that for some permutation $(\frac{c_{1}}{d},\frac{c_{2}}{d},\ldots,\frac{c_{n}}{d})$ of $(\frac{c}{d},\frac{d-c}{d},\frac{1}{d},\ldots,\frac{1}{d})$ we have $\frac{b_{i}-a_{i}}{b_{i}}<\frac{c_{i}}{d}$ for each $i$.
-The following specific examples illustrate that there are many Seifert manifolds $M$ that have co-orientable taut foliations with non-zero Euler class.
-Examples 5.3.
-Using Lemma 5.2, if some $b_{i},b_{j}$ have a common factor greater than $2$, one can construct many examples $M$ with $e(\nu_{M})\neq 0$.
-On the other hand using the conditions (1), (2), (3) immediately above, one can arrange for many of these to have horizontal foliations.
-In (1), (2), (3), and (4) below we give some other classes of examples of Seifert fibred manifolds which admit co-orientable taut foliations but none with vanishing Euler class.
-The proofs are all applications of Theorem 1.1; we omit the details.
-(1) By Lemma 5.2, if $b_{1}=b_{2}=\cdots=b_{n}$ and $e(\nu_{M})=0$, then $a_{1}=a_{2}=\cdots=a_{n}$. The first set of examples deals with a special case of this situation.
-For $n\geq 4$, consider the manifold $M=M(0;b,\frac{1}{c},...,\frac{1}{c})$. Then
-  • (a)
-    $M$ has a horizontal foliation if and only if either $2-n\leq b\leq-2$, or b=1b=-1 and $c\geq 3$.
-  • (b)
-    If $c\geq 2n-2$, then $e(\nu_{M})=0$ if and only if b=2nb=2-n.
-The bound on $c$ in part (b) is optimal: if c=2n3c=2n-3 and b=1b=-1, then $e(\nu_{M})=0$.
-(2) For $n\geq 3$, let $M=M(0;2-n,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$, where $\sum_{i}\frac{a_{i}}{b_{i}}\leq\frac{n-2}{2}$.
-Then $M$ has a horizontal foliation, but $e(\nu_{M})=0$ if and only if $a_{i}=1$ for $1\leq i\leq n$.
-(3) For $n\geq 4$, consider the manifold $M=M(0;b,\frac{1}{b_{1}},\ldots,\frac{1}{b_{n}})$, where the $b_{i}$ are pairwise coprime and $\geq n$. Then
-  • (a)
-    $M$ has a horizontal foliation if and only if $2-n\leq b\leq-1$.
-  • (b)
-    $e(\nu_{M})=0$ if and only if b=2nb=2-n.
-(4) For $n\geq 4$, let $M=M(0;b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$, where $\sum_{i}\frac{a_{i}}{b_{i}}\leq\frac{n}{6}$. Then
-  • (a)
-    $M$ has a horizontal foliation for $2-n\leq b\leq-2$, and for b=1b=-1 if $n=4,5$, or $6$.
-  • (b)
-    $e(\nu_{M})=0$ for at most one $b$ with $2-n\leq b\leq\frac{2-n}{2}$.
+Theorem 5.5.
+Let $\mathcal{B}$ be as above.
+  1. (1)
+    There exists a Seifert fibred rational homology sphere $M$ with base orbifold $\mathcal{B}$ such that $M$ has a horizontal foliation and $e(\nu_{M})=0$.
+  2. (2)
+    With exactly 58 exceptions $\mathcal{B}$, there exists a Seifert fibred rational homology sphere $M$ with base orbifold $\mathcal{B}$ such that $M$ has a horizontal foliation and $e(\nu_{M})\neq 0$.
+Thus the manifolds $M$ in part (2) have co-orientable taut foliations but none with Euler class zero.
+To prove part (1) we take $M=M(0;2-n,1/b_{1},...,1/b_{n})$.
+It follows from Remark 5.4 that $M$ has a horizontal foliation.
+(When $n=3$, use the fact that $\mathcal{B}$ is hyperbolic.) Taking $m=1$ in Theorem 3.3 shows that $e(\nu_{M})=0$.
+The exceptions in part (2) are explicitly listed in the Appendix.
+We will not include a complete proof of part (2); we content ourselves with proving two partial results, Theorems 5.6 and 5.7 below, to illustrate the general strategy.
+First we introduce some notation.
+Let $X=-\chi(\mathcal{B})=(n-2)-\sum_{i=1}^{n}{1/b_{i}}$.
+Since $\mathcal{B}$ is hyperbolic, $X>0$.
+Next let $M$ be a Seifert fibre space with base orbifold $\mathcal{B}$; thus $M$ is of the form $M(0;b,a_{1}/b_{1},...,a_{n}/b_{n})$.
+Then
+$e(M)=-b-\sum_{i=1}^{n}{a_{i}/b_{i}}=X-\left((n-2)+b+\sum_{i=1}^{n}\frac{a_{i}-1}{b_{i}}\right)$
+Let $\alpha=(n-2)+b+\sum_{i=1}^{n}{(a_{i}-1)/b_{i}}$. Then
+$e(M)=X-\alpha$
+By Theorem 3.3, $e(\nu_{M})=0$ if and only if there exists an integer $m$ such that $ma_{i}\equiv 1$ (mod $b_{i}$) for $1\leq i\leq n$ and $X=m(X-\alpha)$, i.e.
+$m\alpha=(m-1)X$.
+We note that if $X-\alpha>0$ then $M$ is a rational homology sphere, and if in addition $\alpha\neq 0$ then $m>1$.
+Theorem 5.6.
+Let $\mathcal{B}=S^{2}(b_{1},...,b_{n})$ be as above, where $n\geq 7$.
+If $n=7$, assume that some $b_{i}$ is greater than $2$.
+Then there is a Seifert fibred rational homology sphere $M$ with base orbifold $\mathcal{B}$ such that $M$ has a horizontal foliation and $e(\nu_{M})\neq 0$.
+Proof.
+We take $M$ to be $M(0;3-n,1/b_{1},...,1/b_{n})$. Then $M$ has a horizontal foliation by Remark 5.4.
+We have $X\geq(n-2)-n/2=(n-4)/2$, with strict inequality if $n=7$. Hence $X>3/2$. Also, $\alpha=1$, so $M$ is a rational homology sphere.
+If $e(\nu_{M})=0$ then there exists an integer $m>1$ such that $m=m\alpha=(m-1)X>(3/2)(m-1)$. Therefore $m<3$, so $m=2$. But then $m\not\equiv 1$ (mod $b_{i}$) for any $i$. ∎
+The exclusion of the case $\mathcal{B}=S^{2}(2,2,...,2)$, $n=7$, is necessary.
+To see this, note that here X=(72)7/2=3/2X=(7-2)-7/2=3/2.
+The manifolds $M$ with base orbifold $\mathcal{B}$ are those of the form M(0;b,1/2,1/2,,1/2)M(0;b,1/2,1/2,...,1/2).
+By Remark 5.4 $M$ has a horizontal foliation if and only if $5\leq b\leq-2$, i.e.
+$\alpha=k$, $0\leq k\leq 3$.
+The corresponding manifold $M_{k}$, say, has $e(\nu_{M_{k}})=0$ if and only if there exists an odd integer $m_{k}$ such that $X=m_{k}(X-k)$.
+Hence $m_{k}=3/(3-2k)$, so $m_{0}=1,m_{1}=3,m_{2}=-3$, and $m_{3}=-1.$ (In fact $M_{3-k}\cong-M_{k}.)$ Thus every $M$ with base orbifold $\mathcal{B}$ that has a horizontal foliation has $e(\nu_{M})=0$, so $\mathcal{B}(2,2,...,2)$, $n=7$, is one of the exceptions in part (2) of Theorem 5.5.
+The next theorem deals with the case $n=3$ when each $b_{i}$ is at least $5$.
+Theorem 5.7.
+Let $\mathcal{B}=S^{2}(b_{1},b_{2},b_{3})$, where $b_{i}\geq 5$, $1\leq i\leq 3$, and the $b_{i}$ are not all $6$.
+Then there is a Seifert fibred rational homology sphere $M$ with base orbifold $\mathcal{B}$ such that $M$ has a horizontal foliation and $e(\nu_{M})\neq 0$.
+The proof will use the following lemma.
+Lemma 5.8.
+If either $c=5$ or $c\geq 7$ then there exists a prime $p$ which does not divide $c$ such that $(p-1)/c\leq 1/3$ and $p/c<1/2$.
+Proof.
+If $c$ is odd take $p=2$.
+If $c$ is even and not a multiple of $3$, take $p=3$.
+So we may suppose that $c=6n$, where $n>1$.
+By Bertrand’s Postulate there exists a prime $p$ such that $n+2<p<2n+2$.
+The left-hand inequality implies that $p$ does not divide $c$ and the right-hand inequality shows that $p-1\leq 2n$.
+Therefore $(p-1)/c\leq 2n/6n=1/3$.
+Also $p/c\leq 1/3+1/c<1/2$.
+∎
+Proof of Theorem 5.7.
+If two of the $b_{i}$, say $b_{1}$ and $b_{2}$, are equal to $5$ then we take $M=M(0;-1,1/5,2/5,1/b_{3})$; see Lemma 5.2. We may therefore assume that X>13/5=2/5.X>1-3/5=2/5.
+By hypothesis, some $b_{i}$, say $b_{1}$, is not equal to $6$.
+Then by Lemma 5.8 there is a prime $p$ such that $(p-1)/b_{1}\leq 1/3$ and $p/b_{1}<1/2$.
+Let $M=M(0;-1,p/b_{1},1/b_{2},1/b_{3})$.
+Since $p/b_{1}<1/2$, $M$ has a horizontal foliation by Remark 5.4.
+Also, $\alpha=(p-1)/b_{1}$, so $X-\alpha>0$.
+Hence $M$ is a rational homology sphere.
+If $e(\nu_{M})=0$ then there exists $m>1$ such that $m\alpha=(m-1)X$.
+Therefore m/3>(2/5)(m1)m/3>(2/5)(m-1).
+Hence 5m>6(m1)5m>6(m-1), so $m<6$.
+But then, since $b_{2}\geq 5$, $m\not\equiv 1$ (mod $b_{2}$).
+∎
+It is necessary to exclude the case $\mathcal{B}=S^{2}(6,6,6)$.
+For it is easy to see from Remark 5.4 that the only manifold $M$ with base orbifold $\mathcal{B}$ that has a horizontal foliation is (up to orientation) M(0;1,1/6,1/6,1/6)M(0;-1,1/6,1/6,1/6), which has $e(\nu_{M})=0$.
+Appendix A Exceptions to Theorem 5.5
+Here we list the orbifolds $S^{2}(b_{1},...,b_{n})$ that are exceptions in part (2) of Theorem 5.5.
+We adopt the notational convention stated at the beginning of §5.3, and also assume that $b_{1}\leq\cdots\leq b_{n}$.
+Let $M=M(0;b,a_{1}/b_{1},...,a_{n}/b_{n})$ be a Seifert fibre space with base orbifold $S^{2}(b_{1},...,b_{n})$. We distinguish three types of exceptional $S^{2}(b_{1},...,b_{n})$.
+  1. (1)
+    $M$ has a horizontal foliation if and only if up to orientation, i.e. for either $M$ or M-M, $a_{1}=\ldots=a_{n}=1$ and b=2nb=2-n. Then taking $m=1$ shows that $e(\nu_{M})=0$.
+  2. (2)
+    There exists $M$ with a horizontal foliation such that for both $M$ and M-M either $b\neq 2-n$ or some $a_{i}\neq 1$, but any such $M$ has $e(\nu_{M})=0$.
+  3. (3)
+    There exists $M$ with a horizontal foliation having $e(\nu_{M})\neq 0$, but any such $M$ has $e(M)=0$, and hence is not a rational homology sphere.
+The exceptions of Type (1) can be easily determined using Remark 5.4. They are
+  • $S^{2}(2,3,c)$, where $c=7,8,9,10,12,14,18,24,$ or $30$
+  • $S^{2}(2,4,c)$, where $c=5,6,8,$ or $12$
+  • $S^{2}(2,5,c)$, where $c=5$ or 66
+  • $S^{2}(2,6,c)$, where $c=6$ or $12$
+  • $S^{2}(3,3,c)$, where $c=4$ or 66
+  • $S^{2}(3,4,c)$, where $c=4$ or 66
+  • $S^{2}(4,4,c)$, where $c=4$ or 66
+  • $S^{2}(6,6,6)$
+  • $S^{2}(2,2,2,c)$, where $c=3,4,$ or 66
+  • $S^{2}(2,2,2,2,2)$
+For the exceptions of Type (2), we list the corresponding manifolds $M$ up to orientation, and the values of $m$ which show that $e(\nu_{M})=0$.
+  • $S^{2}(2,3,c)$, where $c=11,13,16$, or $20$ and M=M(0;1,1/2,1/3,a/c)M=M(0;-1,1/2,1/3,a/c), $1<a<c/2$. In particular,
+    • when $c=11$, $a=2$ and m=5m=-5
+    • when $c=13$, $a=2$ and $m=7$
+    • when $c=16$, $a=3$ and m=5m=-5
+    • when $c=20$, $a=3$ and $m=7$
+  • $S^{2}(2,4,c)$, where $c=7,9,10,14,18,$ or $30$ and M=M(0;1,1/2,1/4,a/c)M=M(0;-1,1/2,1/4,a/c), $1<a<c/2$. In particular,
+    • when $c=7$, $a=2$ and m=3m=-3
+    • when $c=9$, $a=2$ and $m=5$
+    • when $c=10$, $a=3$ and m=3m=-3
+    • when $c=14$, $a=3$ and $m=5$
+    • when $c=18$, $a=3$ and m=7m=-7
+    • when $c=30$, $a=7$ and $m=13$
+  • $S^{2}(2,6,c)$, where $c=8$ or $10$ and M=M(0;1,1/2,1/6,a/c)M=M(0;-1,1/2,1/6,a/c), $1<a<c/2$. In particular,
+    • when $c=8$, $a=3$ and m=5m=-5
+    • when $c=10$, $a=3$ and $m=7$
+  • $S^{2}(3,3,c)$, where $c=5,7,8,$ or $10$ and M=M(0;1,1/3,1/3,a/c)M=M(0;-1,1/3,1/3,a/c), $1<a<c/2$. In particular,
+    • when $c=5$, $a=2$ and m=2m=-2
+    • when $c=7$, $a=2$ and m=5m=-5; or $a=3$ and m=2m=-2
+    • when $c=8$, $a=3$ and m=5m=-5
+    • when $c=10$, $a=3$ and $m=7$
+  • $S^{2}(3,4,5)$, where M=M(0;1,1/3,1/4,2/5)M=M(0;-1,1/3,1/4,2/5) and $m=13$
+  • $S^{2}(4,4,12)$, where M=M(0;1,1/4,1/4,5/12)M=M(0;-1,1/4,1/4,5/12) and $m=5$
+  • $S^{2}(4,6,6)$, where $M=M(3/4,1/6,1/6)$ and m=5m=-5
+  • $S^{2}(2,2,2,c)$, where $c=5,8,$ or $12$ and M=M(0;2,1/2,1/2,1/2,a/c)M=M(0;-2,1/2,1/2,1/2,a/c), $1<a<c/2$. In particular,
+    • when $c=5$, $a=2$ and $m=3$
+    • when $c=8$, $a=3$ and $m=3$
+    • when $c=12$, $a=5$ and $m=5$
+  • $S^{2}(2,2,3,4)$, where M=M(0;2,1/2,1/2,2/3,1/4)M=M(0;-2,1/2,1/2,2/3,1/4) and $m=5$
+  • $S^{2}(2,2,4,6)$, where M=M(0;2,1/2,1/2,3/4,1/6)M=M(0;-2,1/2,1/2,3/4,1/6) and $m=7$
+  • $S^{2}(2,2,2,2,c)$, where $c=4$ or $6$. In particular,
+    • when $c=4$, M=M(0;2,1/2,1/2,1/2,1/2,1/4)M=M(0;-2,1/2,1/2,1/2,1/2,1/4) and m=3m=-3
+    • when $c=6$, M=M(0;2,1/2,1/2,1/2,1/2,1/6)M=M(0;-2,1/2,1/2,1/2,1/2,1/6) and m=5m=-5
+  • $S^{2}(2,2,2,2,2,2,2)$. See §5.3.
+Lastly, we list the exceptions of Type (3).
+In each case, we list the only manifolds $M$, up to orientation, that admit a horizontal foliation with $e(\nu_{M})\neq 0$.
+One easily checks that in all cases $e(M)=0$:
+  • $S^{2}(2,5,10)$, where M=M(0;1,1/2,1/5,3/10)M=M(0;-1,1/2,1/5,3/10) and M=M(0;1,1/2,2/5,1/10)M=M(0;-1,1/2,2/5,1/10)
+  • $S^{2}(3,4,12)$, where M=M(0;1,1/3,1/4,5/12)M=M(0;-1,1/3,1/4,5/12)
+  • $S^{2}(2,2,2,2,2,2)$, where M=M(0;3,1/2,1/2,1/2,1/2,1/2,1/2)M=M(0;-3,1/2,1/2,1/2,1/2,1/2,1/2)